### v1

# Thinking Racial Bias in Fair Forgery Detection: Models, Datasets and Evaluations

Abstract
Due to the successful development of deep image generation technology, forgery detection plays a more important role in social and economic security. Racial bias has not been explored thoroughly in the deep forgery detection field. In the paper, we first contribute a dedicated dataset called the Fair Forgery Detection (FairFD) dataset, where we prove the racial bias of public state-of-the-art (SOTA) methods. Different from existing forgery detection datasets, the self-construct FairFD dataset contains a balanced racial ratio and diverse forgery generation images with the largest-scale subjects. Additionally, we identify the problems with naive fairness metrics when benchmarking forgery detection models. To comprehensively evaluate fairness, we design novel metrics including Approach Averaged Metric and Utility Regularized Metric, which can avoid deceptive results. Extensive experiments conducted with nine representative forgery detection models demonstrate the value of the proposed dataset and the reasonability of the designed fairness metrics. We also conduct more in-depth analyses to offer more insights to inspire researchers in the community.
Keywords Face Forgery Detection  $\cdot$ Racial Bias  $\cdot$ Fairness Evaluation
1 Introduction
Face forgery refers to the creation of fake images or videos of a person’s face using conventional techniques or deep learning methods. These forgeries can be used to spread misinformation, commit fraud, or even blackmail people. There are numerous methods are proposed for detecting face forgery [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Although an increasing number of advanced face forgery detection technologies are being developed, the racial fairness of these detectors is consistently overlooked by researchers [15]. Detectors with severe racial bias can lead to significant social impact. For example, these detectors might disproportionately label faces from a particular racial group as fake, thereby indicating discrimination towards this particular racial group. Therefore, when a detector is ready for deployment, evaluating and analyzing its fairness is a crucial process. However, although there is extensive available research about the fairness in machine learning to draw upon [16, 17, 18, 19, 20], evaluating the fairness in face forgery detection systems remains difficult. This is due to several distinct differences between face forgery detection and other machine learning tasks. The current conditions for evaluating the fairness of face forgery detection systems are still insufficient due to the following reasons:
Existing face forgery detection datasets are not suitable for racial fairness evaluation. We provide the statistical data for current mainstream datasets in Tab 2. And we propose three reasons why these existing datasets are not suitable for racial fairness evaluation. 1.Limited Number of Subjects. Current datasets usually have a limited number of subjects. However, we demonstrate that there is significant fluctuation in the model’s performance across subjects. This result implies these datasets are unsuitable for evaluating racial fairness since individual fairness may overshadow group fairness. Detailed analyses can be found in Sec 3.1. 2.Lack of Diversity in Forgery Approaches. We find that different forgery methods have different fairness levels (see Tab 1). So, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the detectors’ fairness. Detailed description can be found in Sec 3.1. 3.Unbalanced Racial Ratio. Existing face forgery detection datasets are usually not race-balanced (see Tab 2).
Figure 1: Workflow of fairness evaluation in forgery detection. We first construct an evaluation dataset containing a large number of subjects, diverse forgery approaches, and racial balance. Subsequently, we obtain the test results of the forgery detector on each race and forgery method. Finally, we comprehensively evaluate the detector using three sets of 12 fairness metrics in total.
Xception Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.7960 0.7201 0.8449 0.7715 0.7831 0.0450 0.1248
Fake Face FaceSwap 0.8910 0.8949 0.7668 0.8968 0.8624 0.0552 0.1300
SimSwap 0.3950 0.5436 0.2137 0.4264 0.3947 0.1183 0.3299
FaceReen 0.6102 0.6888 0.5223 0.6329 0.6136 0.0599 0.1665
DualReen 0.4114 0.5458 0.2406 0.4691 0.4167 0.1123 0.3052
MaskGAN 0.2593 0.3193 0.2013 0.2626 0.2606 0.0417 0.1180
Overall 0.5605 0.6188 0.4649 0.5766 0.5552 0.0563 0.1538
(a) Xception.
F3Net Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.8117 0.7393 0.7958 0.7695 0.7791 0.0275 0.0724
Fake Face FaceSwap 0.9469 0.9454 0.8665 0.9510 0.9275 0.0352 0.0845
SimSwap 0.3391 0.4590 0.2032 0.3929 0.3486 0.0941 0.2558
FaceReen 0.5992 0.6775 0.6367 0.6773 0.6477 0.0325 0.0783
DualReen 0.1793 0.2838 0.1652 0.2036 0.2080 0.0459 0.1186
MaskGAN 0.2226 0.2643 0.2432 0.2127 0.2357 0.0198 0.0516
Overall 0.5165 0.5616 0.4851 0.5345 0.5244 0.0278 0.0765
(b) F3Net.
RECCE Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.8423 0.7819 0.8799 0.8390 0.8358 0.0350 0.0980
Fake Face FaceSwap 0.8848 0.9036 0.7605 0.8867 0.8589 0.0573 0.1431
SimSwap 0.3054 0.4540 0.1547 0.3489 0.3158 0.1075 0.2993
FaceReen 0.5348 0.6161 0.4667 0.5632 0.5452 0.0539 0.1494
DualReen 0.2298 0.3533 0.1275 0.2583 0.2422 0.0805 0.2258
MaskGAN 0.2002 0.2552 0.1843 0.1947 0.2086 0.0275 0.0709
Overall 0.4996 0.5607 0.4289 0.5151 0.5011 0.0473 0.1318
(c) RECCE.
UCF Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
Caucasian Asian African Indian Mean STD AccGap
Real Face 0.8738 0.8008 0.8984 0.8752 0.8621 0.0367 0.0976
Fake Face FaceSwap 0.9192 0.9314 0.7722 0.9229 0.8864 0.0661 0.1592
SimSwap 0.3558 0.5081 0.1428 0.3488 0.3389 0.1299 0.3653
FaceReen 0.5889 0.7145 0.4831 0.6000 0.5966 0.0819 0.2314
DualReen 0.3756 0.5286 0.2420 0.4091 0.3888 0.1021 0.2866
MaskGAN 0.2615 0.3374 0.2130 0.2326 0.2611 0.0473 0.1244
Overall 0.5625 0.6368 0.4586 0.5648 0.5557 0.0635 0.1782
(d) UCF.
Table 1: Racial bias (STD and AccGap (the maximum differences in accuracy)) evaluations on four popular detectors (Xception [1], F3Net [12], RECCE [10], UCF [11]). The four detectors are trained on FF++ [1] and tested on our FairFD dataset. Training details can be found in Sec 5.1. The results are categorized by ethnicity (Caucasian, Asian, African, Indian). "Mean" donates the average accuracy for four ethnicities. The results are also categorized by real and fake faces. "Overall" represents the average accuracy of the total test dataset. Fake faces are split by forgery approaches.
Existing fairness metrics will cause deceptive results. We identify the bias offset and the aggregation distortion problem, which will cause deceptive racial bias. 1.Bias Offset Problem. Existing fairness metrics typically utilize overall average accuracy for calculations instead of separate calculations for each forgery method. However, this way may obscure part of the bias since different forgery methods may have different privileged races. A detailed explanation of the causes, examples, and solutions for the bias offset is in Sec 4.1. 2.Aggregation Distortion Problem. Current face forgery detectors often suffer from a significant limitation – they struggle to achieve satisfactory performance on unseen forgery methods. As a consequence, these models exhibit varying performance levels across different forgery techniques (See in Tab 1). For existing fairness metrics, even if two forgery methods yield the same bias, detectors with lower baselines are actually more unfair. Treating each forgery method as equal is unreasonable. Detailed explanation is in Sec 4.2.
This work aims to fill the gap in research on racial fairness in face forgery detection by proposing a comprehensive and accurate fairness evaluation system. To achieve this, we introduce new datasets and fairness metrics for assessing existing detectors. We present the workflow of fairness evaluation in Fig 1. The contributions of our work are summarized as follows:
  • $\bullet$
    To our knowledge, we first introduce the forgery detection racial bias evaluation benchmark which comprehensively provides the indicated large-scale dataset, fairness metrics and evaluation protocols.
  • $\bullet$
    We firstly introduce the Fair Forgery Detection (FairFD) dataset for racial bias evaluation, which contains the largest scale subjects, race-balanced ratio and incorporates diverse forgery approaches including FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25].
  • $\bullet$
    We identify the bias offset and aggregation distortion problems with naive fairness metrics. Following, the novel Approach Averaged Metric and Utility Regularized Metric are designed to address the mentioned issues.
  • $\bullet$
    Extensive experimental results prove the limited fairness of SOTA methods and verify the value of the proposed metric. Besides, we also give in-depth analyses and interesting insights to promote the development of the field.
Dataset Race Rate Race Balanced Undefined Ethnicity Subject Number Approaches Real Frame Number Fake Frame Number
Caucasian Asian Indian African Others
FF++ [1] $\sim$43.9$\sim$16.8$\sim$3.2$\sim$3.8$\sim$32.3Yes $\sim$1000 1 73.770k 266.850k
UADFV [26] 97.96% 2.04% 0 0 0 Yes 49 1 241 252
CelebDF-v2 [27] 88.10% 5.10% 0 6.80% 0 Yes 59 1 225.4k 2,116.8k
DFDC [28] - - - - - Yes 960 8 488.4k 1,783.3k
DF-1.0 [29] $\sim$25$\sim$25$\sim$25$\sim$250 Yes 100 1 total 17,600k
ForgeryNet [30] - - - - - Yes 5400 15 1438.201k 1457.861k
FairFD(ours) $\sim$25$\sim$25$\sim$25$\sim$250 No 11430 5 52.041k 253.794k
Table 2: Face Forgery Detection Dataset Comparison. Our dataset is race-balanced, with no undefined races, the number of subjects is maximized, and forgery approach exhibits diversity.
2 Related Work
2.1 Fairness in Face Forgery Detection
The research on fairness in face forgery detection is still relatively limited and waits for further exploration. A preliminary study [31] investigates bias in three commonly used face forgery detectors. They created real face images by sampling from the RFW [32] and UTKFace [33]. Next, they generate fake faces by blending two faces. However, they do not explicitly consider any fairness metrics. This simple study lacks a reasonable evaluation system but still verifies face forgery detectors have a significant racial bias to some extent. Study in [34] annotates five popular deepfake detection datasets with age, gender, ethnicity, etc. Due to the racial imbalance of current datasets, they also propose a metric to deal with the unbalanced test dataset. Another study [35] creat a gender-balanced dataset (GBDF) sampled from the FF++, Celeb-DF, and DF-1.0. This approach involves a limited number of subjects, which restricts the dataset’s capability for fairness evaluations.
Existing research on fairness in face forgery detection is evidently lacking depth. Although a gender-balanced dataset is proposed, it is derived from existing datasets, and our work reveals inherent limitations when using these datasets for evaluation. There is currently no suitable dataset to evaluate the fairness of face forgery detection systems. These limitations are overlooked in previous work. We are the first to conduct in-depth analyses of fairness in face forgery detection and develop a comprehensive and accurate fairness evaluation system.
2.2 Deepfake Datasets and Fairness Metrics
Deepfake Dataset. We summarize the information of existing datasets in Tab 2. We provide the proportions of each race, along with whether the datasets are race-balanced. Additionally, we present whether undefined ethnicity faces are included(,i.e., faces from one race are replaced with another race. We also supply the number of subjects, the number of forgery approaches, and the total number of frames. We also give detailed description of these widely used face forgery datasets in Sec A.3 of Supp.
Fairness Metric. To evaluate racial fairness, what we require is a group fairness metric. There are various group fairness metrics, and the selection of a specific metric also depends on the application context. Commonly, researchers use the performance metric difference between privileged groups and unprivileged groups as a fairness metric. Demographic Parity Difference (DPD) [16, 17] utilizes the difference in positive rate, which represents the proportion of data predicted to be positive, as its fairness metric. The difference in Equalized Odds (DEOdds) [16] utilizes the average of the differences in true positive rate and false positive rate as its fairness metric. The difference in Equal Opportunity (DEO) [18] utilizes the difference in true positive rate solely as a fairness metric. In face recognition scenarios, the Standard Deviation (STD) of performance metrics across different groups is often used [32, 36, 37, 38, 39]. Equity-Scaled Segmentation Performance (ESSP) proposes to evaluate segmentation performance and group fairness simultaneously in medical image segmentation scenarios [19]. In the context of the generative model, the frequency of each group in the generated images is computed, and the average difference in frequency between each pair of groups is calculated as fairness metric [20]. Moreover, there are numerous works proposing more suitable metrics based on the application scenarios. In this work, we consider the commonly used metrics including DPD [16, 17], DEOdds [16], DEO [18], STD [32, 36, 37, 38, 39] and our proposed novel fairness metrics. The definitions of these metrics can be found in Sec A.2 of Supp.
3 FairFD Dataset
Figure 2: Representative examples. We select two images for each ethnicity. The top row represents the original image. Below this, from top to bottom, the forgery methods are FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25].
3.1 Limitations of Current Datasets for Fairness Evaluation
The limitations of existing datasets are as follows:
Limited Number of Subjects. The construction process of the existing face forgery detection datasets involves collecting videos, subsequently creating forgeries, and then segmenting them into frames. As a result, there are typically a small number of subjects and each subject has a large number of frames in these datasets. But in fact, the limited number of subjects makes it challenging to draw meaningful comparisons across groups. Here we do a verification experiment to explain and prove it in Sec A.6 of Supp.
Lack of Diversity in Forgery Approaches. In Tab 2, only DFDC and ForgeryNet employ a variety of forgery techniques. In our research, we find that different forgery methods have different fairness levels. We validate this point in the subsequent Tab 1. So, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the system’s fairness.
Unbalanced Racial Ratio. As stated in [32], a race-balanced test dataset helps ensure a fair evaluation across different race groups. An imbalanced test dataset may result in an unfair evaluation, where the performance of groups with a larger number can exert a more significant impact on fairness evaluation. In Tab 2, only DF-1.0 [29] has a balanced number of ethnic groups. Additionally, in identity-replaced forgery approaches, there is a possibility of faces from one ethnicity being replaced with those from another. In [34], this phenomenon is referred to as "undefined attribute annotation." Undefined attributes can also significantly lower the quality of the evaluation of racial fairness.
3.2 FairFD Description
Having such limitations in existing datasets, we now introduce our dataset, FairFD, aiming to address these shortcomings. FairFD endeavors to overcome previous challenges and provide a more reasonable and comprehensive benchmark for evaluating fairness in face forgery detection. Representative examples of ours are presented in Fig 2. The overview of FairFD can be seen in the Tab 2. Subsequently, we delve into several pivotal facets of our dataset.
Our dataset is an image-level dataset, consisting of original pristine images and 5 kinds of forgery images, i.e., FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25]. Due to face forgery requires using detectors like Dlib [40] to identify faces and some faces are challenging to detect, the number of images of each forgery method is less than the number of pristine images. Specifically, the count of pristine images and the count of images for each forgery method are as follows: Pristine: 52.0k, FaceSwap: 50.9k, SimSwap: 46.7k, FastReen: 52.0k, DualReen: 52.0k, and MaskGAN: 52.0k. In addition to the forgery approach label, our approach also includes labels for four ethnicities (i.e., Caucasian, Asian, African, and Indian). Each ethnicity contains approximately 3000 subjects.
3.3 Source Data Collection
To align with our requirements, which include having racial labels, achieving racial balance, and containing a sufficient number of subjects, we use the Racial Faces in the Wild(RFW) [32] dataset in face recognition filed as pristine images. The RFW dataset comprises face images with four racial labels (,i.e., Caucasian, Asian, African, and Indian), containing approximately 3000 subjects in each racial group, with a roughly equal distribution. Each subject has approximately $3\sim 7$ images. All images in the RFW dataset have a resolution of $400\times 400$ pixels. Besides, the images are carefully selected to maintain similar distributions in terms of age, gender, yaw angle, and pitch angle. RFW meets all our requirements. To reduce the time and human resources required for dataset collection and preprocessing, we directly use RFW as the source data.
3.4 Forgery Approach
To achieve the goal of diversity, we choose various approaches and techniques. We classify the forgery methods into face swap, expression reenactment and attribute manipulation. See details in Sec A.1 of Supp. 1) For face swap, we select FaceSwap [21] and SimSwap [22]. Although both methods result in a face-swapping effect, to ensure approach diversity, we select two distinct face-swapping approaches. FaceSwap [21] is a graphic-based face swap method. Whereas SimSwap [22] is a learning-based face swap method. 2) For expression reenactment, we select FastReen [23] and DualReen [24]. Both face swap and expression reenactment methods involve the use of source and target images to transfer faces or expressions from the target image to the source image. For each subject, we first designate it as the source subject. Next, within the same ethnic group, we randomly select another subject as the target subject, ensuring that the transfer occurs only within a single ethnic group as stated in Sec 3.1. Once the source-target pairs are established, these pairs remain fixed in other identity-replaced forgery approaches. Considering that each subject may have multiple images, for each specific image, we randomly select one image from its appointed target subject for the transfer, ensuring diversity and variability in the dataset. 3) For attribute manipulation, we select MaskGAN [25]. MaskGAN [25] provides an official GUI program that allows manual manipulation of attributes by making use of face parsing. However, as we aim to automate the face forgery process, we randomly select a subset from the nose, glasses, left eye, right eye, left eyebrow, right eyebrow, left ear, right ear, mouth, upper lip, and lower lip. We then apply random dilate and erode operations to the selected parts, enlarging or reducing the chosen regions. Missing parts are filled with skin, resulting in the final manipulated outcome. In summary, we select three types, a total of five forgery approaches, which achieve the diversity requirement.
4 The Proposed Evaluation Metrics
Even though we have obtained a suitable dataset for evaluating deepfake detection’s fairness, we still can not get a credible evaluation result due to existing fairness metrics having two disadvantages. Firstly, the bias offset may lead to an underestimation of racial bias. Secondly, aggregation distortion can result in biased evaluation outcomes favoring specific forgery approaches. Below, we introduce the two disadvantages and their corresponding solutions respectively.
We make corrections to four widely used metrics: DPD [16], DEOdds [16], DEO [18], and STD [32]. For clarity, we leverage DPD to introduce our new metric in the following discussions as an example. The following is the definition of DPD:
$DPD=\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}\,\,|\,\,S=s\right)-P\left(\hat{Y}\,\,|\,\,S=s^{\prime}\right)\right|,$ (1)
where $\hat{Y}$ is the predicted labels. $\mathbb{S}$ represents the set of sensitive attributes, $S\in\mathbb{S}$ and $\mathbb{S}=\{\text{Caucasian},\text{Asian},\text{Indian},\text{African}\}$.
4.1 Bias Offset Problem
Figure 3: A face forgery detection model exhibits different biases for different forgery approaches.
Bias offset refers to bias that will be partially obscured due to the calculation process of existing fairness metrics. Existing fairness metrics do not calculate separately for each forgery method instead of the final averaged performance scores. However, this way may obscure certain biases, which we call bias offset. Taking an example, in Fig 3, face forgery detectors may exhibit different biases for various forgery approaches. We calculate AccGap (the maximum differences in accuracy) and STD (standard deviation) for each forgery approach. In this example, both Forgery Approach 1 (FA1) and Forgery Approach 2 (FA2) exhibit an AccGap greater than 0.2 and an STD greater than 0.07. However, for Forgery Approach 1(FA1), the performance of Caucasians is better than Asians, while for FA2, the performance of Asians is better than Caucasians. In this situation, when calculating the fairness score using the final averaged performance scores, bias is to some extent offset, resulting in a smaller bias score. AccGap is less than 0.2, and the STD is less than 0.07 calculated using average accuracy. We refer to this phenomenon as Bias Offset. In Tab 1, there are some practical examples. Such as in UCF, for forgery approaches DualReen and MaskGAN, the accuracies of Caucasian and Indian will cause a Bias Offset problem. A more reasonable way is to calculate fairness metrics separately for each forgery method and average them. We call this fixed version as Approach Averaged Metric.
$\text{AADPD}=\frac{1}{|\mathbb{F}|}\sum_{f\in\mathbb{F}}{\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}\vphantom{i}}}\left|P\left(\hat{Y}\,|\,S=s,F=f\right)-P\left(\hat{Y}\,|\,S=s^{\prime},F=f\right)\right|,$ (2)
where $\mathbb{F}$ donates real face and forgery approaches. $F\in\mathbb{F}$ and $\mathbb{F}=\{\text{RealFace},$
[…]
4.2 Aggregation Distortion Problem
Various forgery approaches not only exhibit different fairness situations but also demonstrate distinct levels of performance. For example, in Tab 1, we can see that in Xception, the average accuracy for five forgery methods is 86.24%, 39.47%, 61.36%, 41.67%, and 26.06%, respectively. We identify the aggregation distortion problem where even if we calculate fairness scores separately for each forgery method and average them together, the averaged result can achieve a distorted fairness score due to the performance difference.
Directly averaging fairness scores when employing common fairness metrics might lead to misleading conclusions. For instance, consider two approaches with the accuracy of 20% and 80% respectively. We assume that both approaches yield a bias of 10% if we employ DEO as the fairness metric. Then, we calculate a simple average, which is also 10%. This would lead us to focus solely on the absolute differences in error rates without taking into account the variations in baselines. If the racial biases are both calculated to be 10%, the forgery method with only a 20% accuracy would evidently be much more unfair. This oversimplified average fails to capture the substantial disparity in performance between the two methods. To address the aforementioned issue, we propose a fixed version. For each forgery approach, we have:
[…]
where $ACC_{F=f}$ calculates the accuracy of given forgery approach(or real face).
After applying Eq 3 to each forgery method, we calculate their results and then obtain the final fairness score by averaging them. We refer to this approach as Utility Regularized Metric. This nuanced method acknowledges the significance of each forgery approach, providing a more accurate and insightful evaluation of fairness in the context of the diverse fairness and performance landscape. In summary, we recommend not relying on a single fairness metric but rather considering a combination of multiple metrics to collectively reflect the fairness of a detector.
5 Results and Analysis
5.1 Experimental Setup
Dataset. We use FF++ (c23) as our training set. Specifically, for each video, we select 32 frames, crop the facial region, and finally resize it to 256 × 256. We utilize the preprocessed data provided by [41], which has already undergone the aforementioned operations. Our proposed new dataset serves as the testing set. As our dataset inherently consists of face images with backgrounds and bodies removed, there is no need for additional face cropping. Subsequently, we resize the images to 256 × 256 for inference. Note that we still provide the original dataset with a resolution of 400 × 400 for scenarios requiring higher resolution.
Algorithms. We summarize the face forgery detection algorithms in Sec A.4. For a comprehensive and fair analysis, we select several representative algorithms. For spatial-based detectors, we choose Xception [1], RECCE [10], UCF [11], Capsule [5], FFD [8] and CORE [9]. For frequency-based detectors, we select F3Net [12], SPSL [13] and SRM [14]. In detail, these models are trained with the Adam optimization algorithm with a learning rate of 0.0002 and an epoch number of 10. The batch size is 32. And data augmentation methods including image compression, horizontal flip and rotation are applied.
Fairness Metric Spatial-based Frequency-based
Xception RECCE UCF Capsule FFD CORE F3Net SPSL SRM
Naive Metric DPD 0.1538 0.1317 0.1782 0.1428 0.1325 0.0914 0.0764 0.0505 0.1335
DEOdds 0.1672 0.1379 0.1655 0.1267 0.1226 0.0858 0.0877 0.0480 0.0982
DEO 0.2095 0.1777 0.2334 0.1849 0.1684 0.1200 0.1030 0.0567 0.1625
STD 0.0563 0.0473 0.0635 0.0513 0.0470 0.0351 0.0278 0.0211 0.0491
Approach Averaged Metric AADPD 0.1957 0.1644 0.2107 0.1655 0.1579 0.1136 0.1102 0.0592 0.1415
AADEOdds 0.1674 0.1378 0.1655 0.1267 0.1255 0.0887 0.0951 0.0512 0.0984
AADEO 0.2099 0.1777 0.2334 0.1849 0.1742 0.1260 0.1178 0.0633 0.1630
AASTD 0.0721 0.0603 0.0773 0.0613 0.0577 0.0442 0.0425 0.0236 0.0523
Utility Regularized Metric URDPD 0.1314 0.1158 0.1433 0.1013 0.1047 0.0805 0.0769 0.0342 0.0917
URDEOdds 0.1069 0.0908 0.1069 0.0782 0.0802 0.0598 0.0624 0.0315 0.0630
URDEO 0.1437 0.1283 0.1614 0.1128 0.1169 0.0909 0.0842 0.0356 0.1060
URSTD 0.0482 0.0423 0.0523 0.0374 0.0381 0.0312 0.0295 0.0136 0.0338
Table 3: Bias evaluation on FairFD for nine face forgery detectors using Naive Metrics, Approach Averaged Metrics, Utility Regularized Metrics. For each row, we use red to represent the minimum, followed by orange for the second lowest, blue for the third lowest, and dark green for the fourth lowest.
5.2 Benchmarking Fairness of Face Forgery Detectors
The benchmark results (Tab 3) present a comprehensive evaluation of nine face forgery detectors using various fairness metrics. We highlight the four fairest detectors using different colors. Based on the results, we draw the following significant observations:
Current face forgery detectors all exhibit a high degree of racial bias. Even the DPD of SPSL, which has the smallest racial bias, reaches 0.0505. This indicates that the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 5.05%. When separately calculating and averaging for each forgery method, the AADPD reaches 5.92%. On the other hand, for the least fair detector UCF, the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 17.82%. This reminds researchers to address the racial bias in face forgery detection.
Current face forgery detectors exist racial bias variation. Comparing the least fair detector UCF with the most fair detector SPSL, the former’s URDPD is 4.19 times that of the latter, URDEOdds is 3.39 times, URDEO 4.53 times, and URSTD is 3.85 times, showing significant gap in racial bias. Other detectors also exhibit varying degrees of racial bias. Furthermore, we observe that three frequency-based detectors, SPSL, F3Net, and SRM, consistently demonstrate the smallest racial bias across all fairness metrics. This indicates that although current detectors are not explicitly designed for racial fairness, detectors based on different mechanisms still possess varying degrees of racial fairness. We conduct an in-depth investigation into this in Sec 5.3.
5.3 Analyses and Discussions
5.3.1 Balanced Training Dataset.
Data imbalance across races is a prevalent source of bias. To assess the impact of imbalanced training sets on racial bias, we create a dataset that balances across different racial groups. The dataset is obtained by sampling from the FF++. Because FF++ contains a very few number of African and Indian subjects, we can only select 26 subjects from each race. This balanced dataset is employed as the training set. We adopt the same data preprocessing methods and training parameters as the racially imbalanced training set. Notably, due to the reduced dataset size, the number of epochs is doubled to $20$.
Metrics Xception F3Net RECCE UCF
Unbalanced Balanced Unbalanced Balanced Unbalanced Balanced Unbalanced Balanced
Naive Metric DPD 0.1538 0.0852 0.0764 0.0817 0.1317 0.0836 0.1782 0.0908
DEOdds 0.1672 0.0837 0.0877 0.0630 0.1379 0.0663 0.1655 0.0674
DEO 0.2095 0.1076 0.1030 0.1027 0.1777 0.1027 0.2334 0.1104
STD 0.0563 0.0334 0.0278 0.0329 0.0473 0.0314 0.0635 0.0355
Approach Averaged Metric AADPD 0.1957 0.1030 0.1102 0.0996 0.1644 0.1044 0.2107 0.0961
AADEOdds 0.1674 0.0857 0.0951 0.0691 0.1378 0.0746 0.1655 0.0674
AADEO 0.2099 0.1116 0.1178 0.1149 0.1777 0.1193 0.2334 0.1104
AASTD 0.0721 0.0412 0.0425 0.0393 0.0603 0.0405 0.0773 0.0374
Utility Regularized Metric URDPD 0.1314 0.0741 0.0769 0.0770 0.1158 0.0733 0.1433 0.0723
URDEOdds 0.1069 0.0573 0.0624 0.0511 0.0908 0.0505 0.1069 0.0487
URDEO 0.1437 0.0824 0.0842 0.0900 0.1283 0.0847 0.1614 0.0842
URSTD 0.0482 0.0297 0.0295 0.0302 0.0423 0.0284 0.0523 0.0281
Utility Accuracy 0.5552 0.3984 0.5266 0.3308 0.5011 0.4276 0.5557 0.3682
Table 4: Evaluations on the unbalanced and balanced training dataset.
Tab 4 presents the results. For convenience, we put the results of models trained on unbalanced and balanced datasets together. For both Naive Metric and Approach Averaged Metric, we observe that models trained on balanced datasets generally exhibit lower racial bias. Despite the significant improvement brought by a balanced dataset, these detectors still demonstrate a relatively high racial bias. It is crucial to note that, particularly for F3Net, not all metrics demonstrate improvement. These results indicate that training models on a balanced dataset cannot completely address the issue of racial bias, because there are other factors contributing to racial bias. For the Utility Regularized Metric, we note that the enhancement achieved through a balanced training dataset is less significant or may even diminish compared to the Naive Metric and Approach Averaged Metric. This is because this set of metrics is influenced by performance, and the reduced data volume results in a notable model performance drop. This highlights the advantage of the Utility Regularized Metric, i.e., it can reflect the model’s performance.
Based on the aforementioned observations, we argue that utilizing a race-balanced training dataset may not be a recommended way. The fact that deliberately collecting such balanced data in real-world scenarios usually implies discarding a wealth of available imbalanced datasets. Existing datasets are also seldom racially balanced. Furthermore, employing a racially balanced training set does not guarantee effective mitigation of racial bias. In conclusion, training with race-balanced datasets poses challenges, considering the scarcity of such datasets in real-world scenarios and the limited effectiveness in addressing racial bias. Therefore, it is preferable to utilize some other fair learning methods that exhibit better trade-offs.
Figure 4: Probability score distribution of each race.
Xception Utility(Accuracy) ↑ Fairness ↓
Caucasian Asian African Indian STD AccGap
Best Threshold 0.6530 0.7300 0.5380 0.6660 - -
Real Face 0.7960/0.8964 0.7201/0.8813 0.8449/0.8767 0.7715/0.8888 0.0450/0.0075 0.1248/0.0197
FaceSwap 0.8910/0.8265 0.8949/0.7883 0.7668/0.7373 0.8968/0.8264 0.0552/0.0366 0.1300/0.0892
(a) Performance and fairness with 0.5 and optimal thresholds as threshold respectively. Left of ’/’ 0.5. Right of ’/’ is optimal threshold.
Threshold AADPD AADEOdds AADEO AASTD URDPD URDEOdds URDEO URSTD
0.5 0.1274 0.1274 0.1300 0.0501 0.1551 0.1551 0.1507 0.0607
BEST [36] 0.0544 0.0544 0.0892 0.0220 0.0301 0.0301 0.0497 0.0122
(b) Fairness metric results at different thresholds.
Table 5: Threshold Analysis: performance and fairness at different thresholds.
5.3.2 Setting Different Threshold for Each Race.
The study in [36] suggests that in face recognition, the confidence score distributions vary among different race groups. So, they propose setting different thresholds for different races can enhance fairness as well as improve overall performance. We utilize only the real face and FaceSwap portions of FairFD. Firstly, we plot the probability score distribution for each race in Fig 4, revealing distinct differences in the confidence score distribution across different races. The Asian subset shows a higher frequency of high confidence scores, making it more prone to be classified as a fake face. Therefore, a larger threshold can be set for the Asian subset. On the other hand, the African subset exhibits a lower frequency of high probability scores, making it less likely to be classified as a fake face, allowing for a smaller threshold. Next, we set the optimal thresholds, which are values that maximize the overall performance for each racial subset. Tab 5(a) presents the optimal threshold values, as well as the performance and fairness scores when using threshold 0.5 and the optimal thresholds. We observe that, under the condition of maximizing overall performance, there is a certain degree of reduction in racial bias. Tab 5(b) demonstrates racial bias using a fairness metric, showing that indeed, setting different thresholds can reduce racial bias. This conclusion aligns with the findings in [36] consistently.
5.3.3 Analysis of Frequency-Based Detector.
Fairness Metric Xception F3Net RECCE UCF
RGB Grayscale RGB Grayscale RGB Grayscale RGB Grayscale
Naive Metric DPD 0.1538 0.1309 0.0764 0.0780 0.1317 0.0811 0.1782 0.1324
DEOdds 0.1672 0.1439 0.0877 0.0976 0.1379 0.0840 0.1655 0.1512
DEO 0.2095 0.1789 0.1030 0.1035 0.1777 0.1076 0.2334 0.1828
STD 0.0563 0.0470 0.0278 0.0283 0.0473 0.0295 0.0635 0.0475
Approach Averaged Metric AADPD 0.1957 0.1673 0.1102 0.1147 0.1644 0.1097 0.2107 0.1722
AADEOdds 0.1674 0.1439 0.0951 0.1055 0.1378 0.0900 0.1655 0.1512
AADEO 0.2099 0.1789 0.1178 0.1193 0.1777 0.1195 0.2334 0.1828
AASTD 0.0721 0.0614 0.0425 0.0441 0.0603 0.0410 0.0773 0.0658
Utility Regularized Metric URDPD 0.1314 0.1144 0.0769 0.0802 0.1158 0.0785 0.1433 0.1209
URDEOdds 0.1069 0.0935 0.0624 0.0687 0.0908 0.0603 0.1069 0.0986
URDEO 0.1437 0.1249 0.0842 0.0859 0.1283 0.0876 0.1614 0.1321
URSTD 0.0482 0.0419 0.0295 0.0309 0.0423 0.0294 0.0523 0.0461
Table 6: Comparison of fairness on RGB and Grayscale images.
In our previous findings, we observe that frequency-based methods exhibit lower racial bias compared to spatial-based methods. In this section, we provide a preliminary discussion. Spatial-based methods focus on learning forgery clues in the spatial domain, mainly including color mismatch, textures, shapes, and blending boundaries [42]. On the other hand, frequency-based methods concentrate on learning forgery clues in the frequency domain, especially targeting high-frequency information related to blending boundaries, edges, and textures [12]. Frequency-based methods capture less color information, which is also crucial in distinguishing between different racial groups. Consequently, we make the assumption that frequency-based methods’ racial biases are smaller due to these detectors learning less color information. Therefore, we convert the RGB images of our FairFD dataset into grayscale to eliminate color information but retain frequency domain information, and test multiple detectors. The results in Tab 6 demonstrate that the racial bias of spatial-based detectors decreases, while the racial bias of frequency-based methods does not decrease and the changes are small. Thus, color indeed appears to be a contributing factor that leads spatial-based methods to have higher racial bias compared to frequency-based methods. There is still much exploration to be done in the comparative study of frequency and spatial domains, which can significantly contribute to the development of methods for mitigating racial bias. We leave this avenue of research for future investigation.
6 Conclusion
This paper explores how to evaluate the racial fairness of face forgery detectors. We identify numerous disadvantages in existing datasets and fairness metrics, then propose a novel dataset FairFD and two sets of fairness metrics to address these issues. Emphatically, we evaluate the fairness of multiple existing face forgery detectors. The results indicate varying levels of racial bias in current detectors. Further analyses reveal some insights into the emergence of racial bias. In summary, addressing fairness is a crucial progress for the advancement of face forgery detection. We propose several unsolved problems here and hope that our work will help to make progress towards solving them. 1). Although we provide an image-level evaluation dataset, video-level evaluation datasets would still be important. This is because some forgery methods are only applicable at the video level [43, 44, 45], and there are already some detectors based on videos [46, 30] that also suffer from racial bias issues. 2). Although we propose two sets of novel fairness metrics to address the limitations of existing metrics, developing a unified and theoretic metric is still an interesting research direction. 3). Developing fair face forgery detectors that have low racial bias is urgently needed.
Appendix A Appendix
A.1 Classification of Forgery Approaches
There are several approaches for creating fake faces. Here, we provide a brief overview of the classification of forgery methods:
1. Identity-replaced Forgery Approach refers to substituting the original identity in an image, e.g., FaceSwap. FaceSwap [21, 22, 47, 48] involves replacing the face of a person in a video or image with another person’s face. The method usually uses deep learning algorithms to detect and extract the faces of the two people and then swaps them.
2. Identity-remained Forgery Approach retains the original identity and alters other facial attributes, e.g., Attribute Manipulation and Expression Reenactment. Attribute Manipulation [49, 50, 25, 51, 52] involves manipulating the attributes of a face, such as skin color, gender, and eye shape, while preserving the identity of the person. The method usually adopts a GAN. Expression Reenactment [53, 45, 54, 23, 24] involves transferring the facial expressions of one person to another person’s face. The method usually detect and track the facial landmarks of the two people and then transfers the expression from the source face to the target face.
A.2 Existing Fairness Metrics
Due to the complexity of fairness evaluation, numerous fairness metrics are proposed from various perspectives to cater to different scenarios. Therefore, we need to leverage multiple fairness metrics to evaluate the fairness in face forgery detection comprehensively. Here, we present the four most commonly used fairness metrics and outline their respective applicable scenarios.
Demographic Parity Difference (DPD): In the context of face forgery detection, where label 1 represents fake face, DPD reflects the model’s inclination to categorize faces of a specific race as fake. When people perceive the classification of faces as fake as a form of discrimination, we can leverage DPD to assess the extent of bias in the model.
$DPD=\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}\,\,|\,\,S=s\right)-P\left(\hat{Y}\,\,|\,\,S=s^{\prime}\right)\right|,$ (4)
where $\hat{Y}$ is the predicted labels. $\mathbb{S}$ represents the set of sensitive attributes, $S\in\mathbb{S}$ and $\mathbb{S}=\{\text{Caucasian},\text{Asian},\text{Indian},\text{African}\}$.
Difference in Equalized Odds (DEOdds): DPD may fail in certain situations [55]. Considering a scenario where a face forgery detector classifies faces of African and Caucasian individuals as fake at a similar rate, but the model makes different types of errors for the two groups. Specifically, for African faces, the false positive rate is significantly higher than for Caucasian faces, while for Caucasian individuals, the false negative rate is higher. In such a case, we can use DEOdds.
$DEOdds=\frac{1}{2}\underset{y=\left\{0,1\right\}}{\Sigma}\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}|Y=y,S=s\right)-P\left(\hat{Y}|Y=y,S=s^{\prime}\right)\right|.$ (5)
Equal Opportunity (DEO): DEO has more relaxed conditions compared to DEOdds. When assessing the fairness between group A and group B, only the images of faces that are inherently fake need to be considered. DEO solely focuses on true positive rates and does not capture the overall classification differences.
$DEO=\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}|Y=1,S=s\right)-P\left(\hat{Y}|Y=1,S=s^{\prime}\right)\right|.$ (6)
Standard Deviation (STD): STD differs from metrics mentioned above. STD considers the overall variability rather than just the differences between the best and worst-performing ethnicities. We use accuracy as the performance metric.
$STD=std\left(\left\{acc\left(S=s\right)_{s\in\mathbb{S}}\right\}\right),$ (7)
where $std$ is a function that calculates the standard deviation of given list. $acc$ calculates the accuracy of a given race.
A.3 Face Forgery Detection Datasets
We summarize the information of existing datasets in Tab 2. Below we give simple description of some widely used face forgery datasets.
FaceForensics++(FF++) [1] is a forensics dataset that consists of 1000 original video sequences downloaded from the Internet(,i.e.,YouTube). The videos are manipulated with four automated face manipulation methods. For face swap, they use Deepfakes [56] and FaceSwap [21]. For expression reenactment, they use Face2Face [45] and NeuralTextures [57].
UADFV [26] contains videos of varying classes, with each video being classified as either real or fake. The dataset is relatively small, with only 98 videos, but it has been found to be convenient in terms of how the data is formatted.
CelebDF-v2 [27] is a comprehensive collection for deepfake forensics, comprising 590 real videos and 5,639 DeepFake videos featuring celebrities with high-quality. These videos were generated through an enhanced synthesis process, ensuring superior visual quality and better representing the DeepFake content prevalent on the internet.
Deepfake Detection Challenge(DFDC) [28] is created by Facebook in partnership with other industry leaders and academic experts. The dataset consists of 128,154 videos featuring 960 paid actors. The dataset was used in a Kaggle competition to create new and better models to detect manipulated media.
DeeperForensics-1.0(DF-1.0) [29] contains 60,000 videos and 17.6 million frames with 100 consented actors. The actors in DF-1.0 have four skin tones: white, black, yellow, brown, with roughly balanced ratio. Different from other datasets, DF-1.0 attach importance to high-quality and diversity of source face videos. Each actor has various poses, expressions, and illuminations. And DF-1.0 has more real videos than fake videos with a ratio of 5:1.
ForgeryNet [30] is a very large dataset for real-world face forgery detection, with 2.9 million images, 221,247 videos and 15 forgery approaches. And a pipeline of conducting various face forgery approaches are proposed.
A.4 Face Forgery Detection Algorithms Categories
Current face forgery detectors can be roughly divided into two categories: Spatial-based, and Frequency-based. Spatial-based method [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] is based on the spatial domain features(,i.e., forgery clues) of the fake face. Such as RECCE [10] utilizes an encoder to reconstruct real face images, thereby exploring the differences between real and fake images in the spatial domain. The differences in the reconstruction of the two are then used as guidance to train a classifier for detecting deepfakes. Frequency-based method [12, 13, 14] identifies distinctions between real and fake faces in frequency domain, which are used to detect whether the face is forged. Such as F3Net [12] utilizes frequency-aware decomposed image components and local frequency statistics to explore forgery patterns, enabling effective forgery detection.
A.5 Fairness in Face Recognition
An increasing number of researchers are are now paying attention to societal issues of artificial intelligence, including adversarial example [58, 59, 60], privacy protection [61, 62, 63], and fairness concerns.
The research on fairness in face forgery detection is similar to the study of fairness in face recognition. RFW [32] is first introduced as a race balanced test dataset for face recognition. They also propose IMAN which uses a deep information maximization adaptation network to align global distribution to decrease race gap at domain-level, and learns the discriminative target representations at cluster level. Another balanced face recognition dataset BFW [36] is proposed to evaluate the fairness of face recognition system both for gender and ethnic groups and this work also shows variations in the optimal scoring threshold for face-pairs across different subgroups.
The issue of racial bias in machine learning has garnered significant attention in multi fields. The research on fairness in face forgery detection is akin to the study of fairness in face recognition. RFW [32] is first introduced as a balanced test dataset for ethnic group. IMAN [32] uses a deep information maximization adaptation network to align global distribution to decrease race gap at domain-level, and learns the discriminative target representations at cluster level. Another balanced faces dataset BFW [36] is proposed to evaluate the fairness of face recognition system both for gender and ethnic groups and this work also shows variations in the optimal scoring threshold for face-pairs across different subgroups. DebFace [37] uses a de-biasing adversarial network to extract disentangled feature representations for both unbiased face recognition and demographics estimation and adopts adversarial learning to minimize correlation among feature factors so as to abate bias influence. [38] involves multiple preprocessing methods to improve the dual-shot face detector, data re-sampling to balance the data distribution, and multiple data enhancement methods to increase accuracy performance and proposes a linear-combination strategy is adopted to benefit from multi-model fusion. Meta Balanced Network [64] uses meta-learning algorithm to learn adaptive margins in large margin loss to mitigate the algorithmic bias in face recognition models. MixFairFace [39] proposes MixFair Adapter to determine and reduce the identity bias of training samples.
A.6 Number of Subjects is Limited
We propose that the limited number of subjects makes it challenging to draw meaningful comparisons across groups.
Caucasian Asian African Indian
TPR 0.7764 0.7365 0.6289 0.6801
TNR 0.9502 0.9652 0.9314 0.9646
Table 7: TPR and TNR of each race.
Here we do a verification experiment to explain and prove it. We utilize the FF++ dataset, retains only the Caucasian, Asian, African, and Indian subsets. The remaining dataset consists of a total of 677 subjects. For each race, 16 subjects are chosen as the test dataset, and the rest serve as the training dataset (9:1 approximately). We train an Xception model on the training dataset for deepfake detection and evaluate its performance on the test dataset. We follow the hyperparameters specified in [41] and train the model for 10 epochs.
Figure 5: TPR and TNR for each subject.
In Fig 5, we show the TPR and TNR on each subject. We can observe significant fluctuations in the model’s performance across each subject. The presence of extreme results also indicates that subjects significantly influence the model’s classification. Therefore, when the number of subjects is insufficient, it becomes challenging to capture the overall characteristics of a group, making it difficult to draw meaningful comparisons across race groups.
A.7 Visualization in Feature Space
The research on fairness in face forgery detection is similar to the study of fairness in face recognition(see Sec A.5 of Supp for more details). As stated in [32], one reason for racial bias is that different subsets’ features are totally separate. So they take the racial bias as a problem of domain gap and propose IMAN(information maximization adaptation network) to decrease this domain gap. We sample 500 samples for each ethnic group and use the well trained Xception model as the detector to obtain features for these samples. Then, we plot the t-SNE dimensionality reduction graphs for the feature spaces of the four ethnic subsets in Fig 7. Considering that different forgery approaches will cause distinct results, we plot for Real Face, FaceSwap and FaceReen respectively. Unlike in [32], we do not observe distinct separation between different subsets at feature level. Next, we use the MMD(Maximum Mean Discrepancy) to mathematically calculate the feature distances between the Caucasian subset and other ethnic groups. The results in Fig 7 demonstrate that although distances show difference, in comparison to the distances in [32], the distances here are all nearly close to zero. These results are because they consider the face recognition task in [32], so the models tends to learn distinctive features for each subject, resulting in significant differences at the feature level. In contrast, for the task of face forgery detection, the detector does not exhibit the same tendency, leading to smaller differences in feature level. Therefore, attempting to enhance the fairness of face forgery detection from a feature-level perspective may not be feasible.
Figure 6: T-SNE visualization.
Figure 7: MMD.

### v3

@@
 In the paper, we first contribute a dedicated dataset called the Fair Forgery Detection (FairFD) dataset, where we prove the racial bias of public state-of-the-art (SOTA) methods.
-Different from existing forgery detection datasets, the self-construct FairFD dataset contains a balanced racial ratio and diverse forgery generation images with the largest-scale subjects.
+Different from existing forgery detection datasets, the self-constructed FairFD dataset contains a balanced racial ratio and diverse forgery generation images with the largest-scale subjects.
 Additionally, we identify the problems with naive fairness metrics when benchmarking forgery detection models.
 To comprehensively evaluate fairness, we design novel metrics including Approach Averaged Metric and Utility Regularized Metric, which can avoid deceptive results.
-Extensive experiments conducted with nine representative forgery detection models demonstrate the value of the proposed dataset and the reasonability of the designed fairness metrics.
-We also conduct more in-depth analyses to offer more insights to inspire researchers in the community.
-Keywords Face Forgery Detection  $\cdot$ Racial Bias  $\cdot$ Fairness Evaluation
-1 Introduction
+We also present an effective and robust post-processing technique, Bias Pruning with Fair Activations (BPFA), which improves fairness without requiring retraining or weight updates.
+Extensive experiments conducted with 12 representative forgery detection models demonstrate the value of the proposed dataset and the reasonability of the designed fairness metrics.
+By applying the BPFA to the existing fairest detector, we achieve a new SOTA.
+Furthermore, we conduct more in-depth analyses to offer more insights to inspire researchers in the community.
+The source code is available at https://github.com/liudan193/Fairness-Benchmark-for-Face-Forgery-Detection.
+Introduction
 Face forgery refers to the creation of fake images or videos of a person’s face using conventional techniques or deep learning methods.
@@
 Detectors with severe racial bias can lead to significant social impact.
-For example, these detectors might disproportionately label faces from a particular racial group as fake, thereby indicating discrimination towards this particular racial group.
+These detectors might disproportionately label faces from a particular racial group as fake, thereby indicating discrimination towards this particular racial group.
 Therefore, when a detector is ready for deployment, evaluating and analyzing its fairness is a crucial process.
 However, although there is extensive available research about the fairness in machine learning to draw upon [16, 17, 18, 19, 20], evaluating the fairness in face forgery detection systems remains difficult.
-This is due to several distinct differences between face forgery detection and other machine learning tasks.
-The current conditions for evaluating the fairness of face forgery detection systems are still insufficient due to the following reasons:
-Existing face forgery detection datasets are not suitable for racial fairness evaluation.
-We provide the statistical data for current mainstream datasets in Tab 2.
-And we propose three reasons why these existing datasets are not suitable for racial fairness evaluation.
-1.Limited Number of Subjects.
-Current datasets usually have a limited number of subjects.
-However, we demonstrate that there is significant fluctuation in the model’s performance across subjects.
-This result implies these datasets are unsuitable for evaluating racial fairness since individual fairness may overshadow group fairness.
-Detailed analyses can be found in Sec 3.1.
-2.Lack of Diversity in Forgery Approaches.
-We find that different forgery methods have different fairness levels (see Tab 1).
-So, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the detectors’ fairness.
-Detailed description can be found in Sec 3.1.
-3.Unbalanced Racial Ratio.
-Existing face forgery detection datasets are usually not race-balanced (see Tab 2).
+This is due to several distinct differences between face forgery detection and other deep learning tasks.
+This work aims to fill the gap in research on racial fairness in face forgery detection by proposing an accurate, comprehensive and credible fairness evaluation system.
+To achieve this goal, we analyze the shortcomings of existing evaluation components (i.e.
+dataset and metric), and our corresponding solutions.
+(1) Dataset.
+Existing face forgery detection datasets have a limited number of subjects.
+We find performance fluctuations significantly across subjects, so individual fairness may overshadow group fairness, which will make the evaluation results inaccurate.
+It also can be found that different forgery approaches have different fairness levels, limited forgery approaches will lead to a non-comprehensive result.
+Otherwise, undefined ethnicity (faces from two ethnicities are swapped) will lead to an inaccurate result.
+(2) Metric.
+We also propose two issues (Bias Offset and Aggregation Distortion) that will cause deceptive results.
+Bias Offset arises because existing fairness metrics typically use overall average accuracy for calculations instead of assessing each forgery method separately.
+This way may obscure some biases as different forgery methods may have different privileged races.
+Aggregation Distortion arises because detectors often show significant utility variations across different forgery techniques.
+Even if two forgery methods exhibit the same bias, detectors with lower utility can be more unfair.
+Treating each forgery method as equal will lead to unreliable results.
 Figure 1: Workflow of fairness evaluation in forgery detection.
@@
 Finally, we comprehensively evaluate the detector using three sets of 12 fairness metrics in total.
-Xception Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
-Caucasian Asian African Indian Mean STD AccGap
-Real Face 0.7960 0.7201 0.8449 0.7715 0.7831 0.0450 0.1248
-Fake Face FaceSwap 0.8910 0.8949 0.7668 0.8968 0.8624 0.0552 0.1300
-SimSwap 0.3950 0.5436 0.2137 0.4264 0.3947 0.1183 0.3299
-FaceReen 0.6102 0.6888 0.5223 0.6329 0.6136 0.0599 0.1665
-DualReen 0.4114 0.5458 0.2406 0.4691 0.4167 0.1123 0.3052
-MaskGAN 0.2593 0.3193 0.2013 0.2626 0.2606 0.0417 0.1180
-Overall 0.5605 0.6188 0.4649 0.5766 0.5552 0.0563 0.1538
-(a) Xception.
-F3Net Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
-Caucasian Asian African Indian Mean STD AccGap
-Real Face 0.8117 0.7393 0.7958 0.7695 0.7791 0.0275 0.0724
-Fake Face FaceSwap 0.9469 0.9454 0.8665 0.9510 0.9275 0.0352 0.0845
-SimSwap 0.3391 0.4590 0.2032 0.3929 0.3486 0.0941 0.2558
-FaceReen 0.5992 0.6775 0.6367 0.6773 0.6477 0.0325 0.0783
-DualReen 0.1793 0.2838 0.1652 0.2036 0.2080 0.0459 0.1186
-MaskGAN 0.2226 0.2643 0.2432 0.2127 0.2357 0.0198 0.0516
-Overall 0.5165 0.5616 0.4851 0.5345 0.5244 0.0278 0.0765
-(b) F3Net.
-RECCE Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
-Caucasian Asian African Indian Mean STD AccGap
-Real Face 0.8423 0.7819 0.8799 0.8390 0.8358 0.0350 0.0980
-Fake Face FaceSwap 0.8848 0.9036 0.7605 0.8867 0.8589 0.0573 0.1431
-SimSwap 0.3054 0.4540 0.1547 0.3489 0.3158 0.1075 0.2993
-FaceReen 0.5348 0.6161 0.4667 0.5632 0.5452 0.0539 0.1494
-DualReen 0.2298 0.3533 0.1275 0.2583 0.2422 0.0805 0.2258
-MaskGAN 0.2002 0.2552 0.1843 0.1947 0.2086 0.0275 0.0709
-Overall 0.4996 0.5607 0.4289 0.5151 0.5011 0.0473 0.1318
-(c) RECCE.
-UCF Utility(Accuracy) $\uparrow$ Fairness $\downarrow$
-Caucasian Asian African Indian Mean STD AccGap
-Real Face 0.8738 0.8008 0.8984 0.8752 0.8621 0.0367 0.0976
-Fake Face FaceSwap 0.9192 0.9314 0.7722 0.9229 0.8864 0.0661 0.1592
-SimSwap 0.3558 0.5081 0.1428 0.3488 0.3389 0.1299 0.3653
-FaceReen 0.5889 0.7145 0.4831 0.6000 0.5966 0.0819 0.2314
-DualReen 0.3756 0.5286 0.2420 0.4091 0.3888 0.1021 0.2866
-MaskGAN 0.2615 0.3374 0.2130 0.2326 0.2611 0.0473 0.1244
-Overall 0.5625 0.6368 0.4586 0.5648 0.5557 0.0635 0.1782
-(d) UCF.
-Table 1: Racial bias (STD and AccGap (the maximum differences in accuracy)) evaluations on four popular detectors (Xception [1], F3Net [12], RECCE [10], UCF [11]).
-The four detectors are trained on FF++ [1] and tested on our FairFD dataset.
-Training details can be found in Sec 5.1.
-The results are categorized by ethnicity (Caucasian, Asian, African, Indian).
-"Mean" donates the average accuracy for four ethnicities.
-The results are also categorized by real and fake faces.
-"Overall" represents the average accuracy of the total test dataset.
-Fake faces are split by forgery approaches.
-Existing fairness metrics will cause deceptive results.
-We identify the bias offset and the aggregation distortion problem, which will cause deceptive racial bias.
-1.Bias Offset Problem.
-Existing fairness metrics typically utilize overall average accuracy for calculations instead of separate calculations for each forgery method.
-However, this way may obscure part of the bias since different forgery methods may have different privileged races.
-A detailed explanation of the causes, examples, and solutions for the bias offset is in Sec 4.1.
-2.Aggregation Distortion Problem.
-Current face forgery detectors often suffer from a significant limitation – they struggle to achieve satisfactory performance on unseen forgery methods.
-As a consequence, these models exhibit varying performance levels across different forgery techniques (See in Tab 1).
-For existing fairness metrics, even if two forgery methods yield the same bias, detectors with lower baselines are actually more unfair.
-Treating each forgery method as equal is unreasonable.
-Detailed explanation is in Sec 4.2.
-This work aims to fill the gap in research on racial fairness in face forgery detection by proposing a comprehensive and accurate fairness evaluation system.
-To achieve this, we introduce new datasets and fairness metrics for assessing existing detectors.
-We present the workflow of fairness evaluation in Fig 1.
-The contributions of our work are summarized as follows:
+To tackle these problems, we firstly introduce the Fair Forgery Detection (FairFD) dataset for racial bias evaluation, which contains the largest scale subjects, and incorporates diverse forgery approaches including Face Swapping: FaceSwap [21], SimSwap [22], Expression Reenactment: FastReen [23], DualReen [24], Face Editing: MaskGAN [25], StarGAN [26], StyleGAN [27], Diffusion-Based: SDSwap [28], DCFace [29], Face2Diffusion [30] and Transformer-Based: FSRT [31].
+And the self-constructed FairFD dataset does not have any undefined ethnicity annotations.
+For the specific metric, we address the mentioned two issues by introducing the Approach Averaged Metric, which calculates fairness separately for each forgery approach and then aggregates them, and the Utility Regularized Metric, which uses the utility to regularize the fairness.
+In addition to our evaluation system, we also introduce a new pruning approach called BPFA (Bias Pruning with Fair Activations).
+The designed BPFA leverages an innovative pruning metric to identify weights with the least impact on model utility while contributing most significantly to bias (e.g., racial bias).
+By pruning these weights, BPFA successfully enhances fairness without compromising utility.
+As a post-processing method, BPFA enhances fairness without any retraining.
+And it can be applied to any detector, including those already trained with existing fairness learning strategies, to further improve fairness performance.
+The workflow of fairness evaluation is illustrated in Figure 1.
+Sufficient experimental results prove that the proposed BPFA is an efficient, plug-and-play and robust pruning scheme, outperforming other baseline pruning methods by a significant margin.
+The key contributions are summarized as follows:
   • $\bullet$
-    To our knowledge, we first introduce the forgery detection racial bias evaluation benchmark which comprehensively provides the indicated large-scale dataset, fairness metrics and evaluation protocols.
-  • $\bullet$
-    We firstly introduce the Fair Forgery Detection (FairFD) dataset for racial bias evaluation, which contains the largest scale subjects, race-balanced ratio and incorporates diverse forgery approaches including FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25].
+    To our knowledge, it is the early exploration to introduce a comprehensive racial bias evaluation benchmark for forgery detection, providing a large-scale dataset, fairness metrics, and evaluation protocols.
+We newly introduce the Fair Forgery Detection (FairFD) dataset for racial bias in forgery detection evaluation, which contains the largest scale subjects, race-balanced ratio and incorporates diverse forgery approaches.
   • $\bullet$
@@
 Following, the novel Approach Averaged Metric and Utility Regularized Metric are designed to address the mentioned issues.
+Extensive experimental results demonstrate the limited fairness of existing SOTA methods and validate the value of our proposed metric.
   • $\bullet$
-    Extensive experimental results prove the limited fairness of SOTA methods and verify the value of the proposed metric.
-Besides, we also give in-depth analyses and interesting insights to promote the development of the field.
-Dataset Race Rate Race Balanced Undefined Ethnicity Subject Number Approaches Real Frame Number Fake Frame Number
+    We propose the Bias Pruning with Fair Activations (BPFA) method to improve the fairness of forgery detectors.
+Extensive experiments demonstrate the advantages of our method, particularly its efficiency, plug-and-play nature and robustness.
+By combining BPFA with the fairest detector, we achieve a new SOTA in racial fairness performance.
+Specifically, we offer in-depth analyses and insightful observations to advance the community.
+Dataset Race Rate Race Balance Undefined Ethnicity Subject Number App- roach Real Img Number Fake Img Number
 Caucasian Asian Indian African Others
-FF++ [1] $\sim$43.9$\sim$16.8$\sim$3.2$\sim$3.8$\sim$32.3Yes $\sim$1000 1 73.770k 266.850k
+FF++ [1] $\sim$43.9$\sim$16.8$\sim$3.2$\sim$3.8$\sim$32.3Yes $\sim$1000 1 73k 266k
 UADFV [32] 97.96% 2.04% 0 0 0 Yes 49 1 241 252
-CelebDF-v2 [27] 88.10% 5.10% 0 6.80% 0 Yes 59 1 225.4k 2,116.8k
-DFDC [28] - - - - - Yes 960 8 488.4k 1,783.3k
+CelebDF-v2 [33] 88.10% 5.10% 0 6.80% 0 Yes 59 1 225k 2,116k
+DFDC [34] - - - - - Yes 960 8 488k 1,783k
 DF-1.0 [35] $\sim$25$\sim$25$\sim$25$\sim$250 Yes 100 1 total 17,600k
-ForgeryNet [30] - - - - - Yes 5400 15 1438.201k 1457.861k
-FairFD(ours) $\sim$25$\sim$25$\sim$25$\sim$250 No 11430 5 52.041k 253.794k
-Table 2: Face Forgery Detection Dataset Comparison. Our dataset is race-balanced, with no undefined races, the number of subjects is maximized, and forgery approach exhibits diversity.
-2 Related Work
-2.1 Fairness in Face Forgery Detection
-The research on fairness in face forgery detection is still relatively limited and waits for further exploration.
-A preliminary study [31] investigates bias in three commonly used face forgery detectors.
-They created real face images by sampling from the RFW [32] and UTKFace [33].
-Next, they generate fake faces by blending two faces.
-However, they do not explicitly consider any fairness metrics.
-This simple study lacks a reasonable evaluation system but still verifies face forgery detectors have a significant racial bias to some extent.
-Study in [34] annotates five popular deepfake detection datasets with age, gender, ethnicity, etc.
-Due to the racial imbalance of current datasets, they also propose a metric to deal with the unbalanced test dataset.
-Another study [35] creat a gender-balanced dataset (GBDF) sampled from the FF++, Celeb-DF, and DF-1.0.
-This approach involves a limited number of subjects, which restricts the dataset’s capability for fairness evaluations.
-Existing research on fairness in face forgery detection is evidently lacking depth.
-Although a gender-balanced dataset is proposed, it is derived from existing datasets, and our work reveals inherent limitations when using these datasets for evaluation.
-There is currently no suitable dataset to evaluate the fairness of face forgery detection systems.
-These limitations are overlooked in previous work.
-We are the first to conduct in-depth analyses of fairness in face forgery detection and develop a comprehensive and accurate fairness evaluation system.
-2.2 Deepfake Datasets and Fairness Metrics
+ForgeryNet [36] - - - - - Yes 5400 15 1438k 1457k
+FairFD(ours) $\sim$25$\sim$25$\sim$25$\sim$250 No 11430 11 52k 572k
+Table 1: Face Forgery Detection Dataset Comparison. Our dataset is race-balanced, with no undefined races, the maximized number of subjects and forgery approaches exhibit diversity.
+Related Work
+Fairness in Face Forgery Detection
+Fairness Algorithm.
+Fairness in face forgery detection is a relatively novel topic.
+DAG-FDD [37] is first proposed to address fairness without demographic information by setting a probability threshold for minority groups to ensure low error rates for all groups meeting this threshold.
+DAW-FDD [37] utilizes demographic information to design losses to ensure similar performance across specified groups.
+PFGDFD [38] improves fairness by using disentanglement loss to separate demographic and forgery features.
 Deepfake Dataset.
-We summarize the information of existing datasets in Tab 2.
+We summarize the information of existing datasets in Table 1.
 We provide the proportions of each race, along with whether the datasets are race-balanced.
@@
 We also supply the number of subjects, the number of forgery approaches, and the total number of frames.
-We also give detailed description of these widely used face forgery datasets in Sec A.3 of Supp.
+DAG(W)-FDD [37] and PFGDFD [38] directly use several of the datasets mentioned above as test data.
+Our work reveals inherent limitations when using these datasets for evaluation.
+There is currently no suitable dataset to evaluate the fairness of forgery detection.
+We also give details of widely used face forgery datasets in the section "Face Forgery Detection Datasets" in Supp.
 Fairness Metric.
-To evaluate racial fairness, what we require is a group fairness metric.
-There are various group fairness metrics, and the selection of a specific metric also depends on the application context.
-Commonly, researchers use the performance metric difference between privileged groups and unprivileged groups as a fairness metric.
-Demographic Parity Difference (DPD) [16, 17] utilizes the difference in positive rate, which represents the proportion of data predicted to be positive, as its fairness metric.
-The difference in Equalized Odds (DEOdds) [16] utilizes the average of the differences in true positive rate and false positive rate as its fairness metric.
-The difference in Equal Opportunity (DEO) [18] utilizes the difference in true positive rate solely as a fairness metric.
-In face recognition scenarios, the Standard Deviation (STD) of performance metrics across different groups is often used [32, 36, 37, 38, 39].
-Equity-Scaled Segmentation Performance (ESSP) proposes to evaluate segmentation performance and group fairness simultaneously in medical image segmentation scenarios [19].
-In the context of the generative model, the frequency of each group in the generated images is computed, and the average difference in frequency between each pair of groups is calculated as fairness metric [20].
-Moreover, there are numerous works proposing more suitable metrics based on the application scenarios.
+To evaluate racial fairness, what we require is group fairness metrics.
+There are various group fairness metrics, and the selection of a metric depends on the application context.
 In this work, we consider the commonly used metrics including DPD [16, 17], DEOdds [16], DEO [18], STD [39, 40, 41, 42, 43] and our proposed novel fairness metrics.
-The definitions of these metrics can be found in Sec A.2 of Supp.
-3 FairFD Dataset
-Figure 2: Representative examples.
-We select two images for each ethnicity.
-The top row represents the original image.
-Below this, from top to bottom, the forgery methods are FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25].
-3.1 Limitations of Current Datasets for Fairness Evaluation
-The limitations of existing datasets are as follows:
+The definitions of these metrics can be found in the section "Existing Fairness Metrics" in Supp.
+FairFD Dataset
+Limitations of Current Datasets
 Limited Number of Subjects.
-The construction process of the existing face forgery detection datasets involves collecting videos, subsequently creating forgeries, and then segmenting them into frames.
+The construction process of the existing face forgery detection datasets involves collecting videos, subsequently creating forgeries, and then extracting frames.
 As a result, there are typically a small number of subjects and each subject has a large number of frames in these datasets.
-But in fact, the limited number of subjects makes it challenging to draw meaningful comparisons across groups.
-Here we do a verification experiment to explain and prove it in Sec A.6 of Supp.
-Lack of Diversity in Forgery Approaches.
-In Tab 2, only DFDC and ForgeryNet employ a variety of forgery techniques.
-In our research, we find that different forgery methods have different fairness levels.
-We validate this point in the subsequent Tab 1.
-So, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the system’s fairness.
-Unbalanced Racial Ratio.
-As stated in [32], a race-balanced test dataset helps ensure a fair evaluation across different race groups.
-An imbalanced test dataset may result in an unfair evaluation, where the performance of groups with a larger number can exert a more significant impact on fairness evaluation.
-In Tab 2, only DF-1.0 [29] has a balanced number of ethnic groups.
-Additionally, in identity-replaced forgery approaches, there is a possibility of faces from one ethnicity being replaced with those from another.
-In [34], this phenomenon is referred to as "undefined attribute annotation." Undefined attributes can also significantly lower the quality of the evaluation of racial fairness.
-3.2 FairFD Description
-Having such limitations in existing datasets, we now introduce our dataset, FairFD, aiming to address these shortcomings.
-FairFD endeavors to overcome previous challenges and provide a more reasonable and comprehensive benchmark for evaluating fairness in face forgery detection.
-Representative examples of ours are presented in Fig 2.
-The overview of FairFD can be seen in the Tab 2.
+The limited number of subjects makes it challenging to draw meaningful comparisons across groups.
+Furthermore, we conducted the verification experiment to analyze and prove it in the section "Number of Subjects is Limited" in Supp.
+Lack of Diversity of Forgery Approaches.
+In Table 1, only DFDC and ForgeryNet employ a variety of forgery techniques.
+However, we find that different forgery methods have different fairness levels.
+We validate this point in the subsequent Figure 4.
+Thus, we should strive to diversify forgery methods as much as possible, enabling a more comprehensive evaluation of the system’s fairness.
+Undefined Attribute Annotation.
+For these identity-replaced forgery approaches, there is a possibility of faces from one ethnicity being replaced with those from another.
+In related work  [44], this phenomenon is referred to as "undefined attribute annotation." Undefined attributes can also significantly lower the quality of the evaluation of racial fairness, which is ignored in existing face forgery detection datasets.
+FairFD Description
+Considering the mentioned limitations in existing datasets, we introduce our dataset, FairFD, aiming to address these shortcomings.
+FairFD endeavors to overcome previous challenges and provide a more accurate, reliable and comprehensive benchmark for evaluating fairness in face forgery detection.
+Representative examples of ours are presented in Figure 1.
+The overview of FairFD can be shown in the Table 1.
 Subsequently, we delve into several pivotal facets of our dataset.
-Our dataset is an image-level dataset, consisting of original pristine images and 5 kinds of forgery images, i.e., FaceSwap [21], SimSwap [22], FastReen [23], DualReen [24], and MaskGAN [25].
-Due to face forgery requires using detectors like Dlib [40] to identify faces and some faces are challenging to detect, the number of images of each forgery method is less than the number of pristine images.
-Specifically, the count of pristine images and the count of images for each forgery method are as follows: Pristine: 52.0k, FaceSwap: 50.9k, SimSwap: 46.7k, FastReen: 52.0k, DualReen: 52.0k, and MaskGAN: 52.0k.
+Our dataset is an image-level dataset, and for each image, there are 11 kinds of corresponding forgery images, i.e., Face Swapping: FaceSwap [21], SimSwap [22], Expression Reenactment: FastReen [23], DualReen [24], Face Editing: StarGAN [26], StyleGAN [27], MaskGAN [25], Diffusion-Based: SDSwap [28], DCFace [29], Face2Diffusion [30] and Transformer-Based: FSRT [31].
 In addition to the forgery approach label, our approach also includes labels for four ethnicities (i.e., Caucasian, Asian, African, and Indian).
 Each ethnicity contains approximately 3000 subjects.
-3.3 Source Data Collection
-To align with our requirements, which include having racial labels, achieving racial balance, and containing a sufficient number of subjects, we use the Racial Faces in the Wild(RFW) [32] dataset in face recognition filed as pristine images.
-The RFW dataset comprises face images with four racial labels (,i.e., Caucasian, Asian, African, and Indian), containing approximately 3000 subjects in each racial group, with a roughly equal distribution.
+Source Data Collection and Forgery Process
+To align with our requirements, which include having racial labels, and containing a sufficient number of subjects, we use the RFW [39] dataset as pristine images.
+The RFW dataset comprises face images with four racial labels (,i.e., Caucasian, Asian, African, and Indian), containing approximately 3000 subjects for each racial group, with a roughly equal distribution.
 Each subject has approximately $3\sim 7$ images.
@@
 Besides, the images are carefully selected to maintain similar distributions in terms of age, gender, yaw angle, and pitch angle.
-RFW meets all our requirements.
-To reduce the time and human resources required for dataset collection and preprocessing, we directly use RFW as the source data.
-3.4 Forgery Approach
+See details in "Detailed Distribution of FairFD" in Supp.
+To reduce the human resources, we use RFW as the source data.
 To achieve the goal of diversity, we choose various approaches and techniques.
-We classify the forgery methods into face swap, expression reenactment and attribute manipulation.
-See details in Sec A.1 of Supp.
-1) For face swap, we select FaceSwap [21] and SimSwap [22].
-Although both methods result in a face-swapping effect, to ensure approach diversity, we select two distinct face-swapping approaches.
-FaceSwap [21] is a graphic-based face swap method.
-Whereas SimSwap [22] is a learning-based face swap method.
-2) For expression reenactment, we select FastReen [23] and DualReen [24].
-Both face swap and expression reenactment methods involve the use of source and target images to transfer faces or expressions from the target image to the source image.
-For each subject, we first designate it as the source subject.
-Next, within the same ethnic group, we randomly select another subject as the target subject, ensuring that the transfer occurs only within a single ethnic group as stated in Sec 3.1.
-Once the source-target pairs are established, these pairs remain fixed in other identity-replaced forgery approaches.
-Considering that each subject may have multiple images, for each specific image, we randomly select one image from its appointed target subject for the transfer, ensuring diversity and variability in the dataset.
-3) For attribute manipulation, we select MaskGAN [25].
-MaskGAN [25] provides an official GUI program that allows manual manipulation of attributes by making use of face parsing.
-However, as we aim to automate the face forgery process, we randomly select a subset from the nose, glasses, left eye, right eye, left eyebrow, right eyebrow, left ear, right ear, mouth, upper lip, and lower lip.
-We then apply random dilate and erode operations to the selected parts, enlarging or reducing the chosen regions.
-Missing parts are filled with skin, resulting in the final manipulated outcome.
-In summary, we select three types, a total of five forgery approaches, which achieve the diversity requirement.
-4 The Proposed Evaluation Metrics
-Even though we have obtained a suitable dataset for evaluating deepfake detection’s fairness, we still can not get a credible evaluation result due to existing fairness metrics having two disadvantages.
+We classify the forgery methods into face swap, expression reenactment, attribute manipulation and advanced forgery methods(stable diffusion and transformer).
+See details in the section "Classification of Forgery Approaches" in Supp.
+We reimplement these methods and apply them to the source data.
+Details configuration and process of forgery crafting can be found in the section "Forgery Crafting Process" in Supp.
+The Proposed Evaluation Metrics
+Even though we have obtained a reliable dataset for evaluating deepfake detection’s fairness, we still can not get a credible evaluation result due to existing fairness metrics having two flaws.
 Firstly, the bias offset may lead to an underestimation of racial bias.
 Secondly, aggregation distortion can result in biased evaluation outcomes favoring specific forgery approaches.
-Below, we introduce the two disadvantages and their corresponding solutions respectively.
+Below, we introduce the two flaws and their corresponding solutions respectively.
 We make corrections to four widely used metrics: DPD [16], DEOdds [16], DEO [18], and STD [39].
@@
 where $\hat{Y}$ is the predicted labels. $\mathbb{S}$ represents the set of sensitive attributes, ${s\in\mathbb{S}}$ and $\mathbb{S}=\{\text{Caucasian},\text{Asian},\text{Indian},\text{African}\}$.
-4.1 Bias Offset Problem
+Bias Offset Problem
 Figure 2: A face forgery detection model exhibits different biases for different forgery approaches.
@@
 However, this way may obscure certain biases, which we call bias offset.
-Taking an example, in Fig 3, face forgery detectors may exhibit different biases for various forgery approaches.
+Taking an example, in Figure 2, face forgery detectors may exhibit different biases for various forgery approaches.
 We calculate AccGap (the maximum differences in accuracy) and STD (standard deviation) for each forgery approach.
@@
 We refer to this phenomenon as bias offset.
-In Tab 1, there are some practical examples.
-Such as in UCF, for forgery approaches DualReen and MaskGAN, the accuracies of Caucasian and Indian will cause a Bias Offset problem.
-A more reasonable way is to calculate fairness metrics separately for each forgery method and average them.
-We call this fixed version as Approach Averaged Metric.
-$\text{AADPD}=\frac{1}{|\mathbb{F}|}\sum_{f\in\mathbb{F}}{\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}\vphantom{i}}}\left|P\left(\hat{Y}\,|\,S=s,F=f\right)-P\left(\hat{Y}\,|\,S=s^{\prime},F=f\right)\right|,$ (2)
-where $\mathbb{F}$ donates real face and forgery approaches. $F\in\mathbb{F}$ and $\mathbb{F}=\{\text{RealFace},$
-[…]
-4.2 Aggregation Distortion Problem
+A more reliable way is to calculate fairness metrics separately for each forgery method and average them.
+We call this novel strategy as Approach Averaged Metric.
+AADPD $=\frac{1}{|\mathbb{F}|}\sum_{f\in\mathbb{F}}{\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}\vphantom{i}}}\bigl{|}P\left(\hat{Y}\,|\,S=s,F=f\right)$
+$-P\left(\hat{Y}\,|\,S=s^{\prime},F=f\right)\bigr{|},$ (2)
+where $\mathbb{F}$ donates real face and forgery approaches. $f\in\mathbb{F}$ and $\mathbb{F}$ is the set of forgery methods.
+Aggregation Distortion Problem
 Various forgery approaches not only exhibit different fairness situations but also demonstrate distinct levels of performance.
-For example, in Tab 1, we can see that in Xception, the average accuracy for five forgery methods is 86.24%, 39.47%, 61.36%, 41.67%, and 26.06%, respectively.
+For example, in Figure.
+3 and Figure.
+4, this is clearly evident.
 We identify the aggregation distortion problem where even if we calculate fairness scores separately for each forgery method and average them together, the averaged result can achieve a distorted fairness score due to the performance difference.
@@
 This oversimplified average fails to capture the substantial disparity in performance between the two methods.
-To address the aforementioned issue, we propose a fixed version.
+To address this issue, we propose a fixed version.
 For each forgery approach, we have:
-[…]
-where $ACC_{F=f}$ calculates the accuracy of given forgery approach(or real face).
-After applying Eq 3 to each forgery method, we calculate their results and then obtain the final fairness score by averaging them.
+URDPD $=\frac{1}{|\mathbb{F}|}\sum_{f\in\mathbb{F}}{\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}\vphantom{i}}}\bigl{|}P\left(\hat{Y}\,|\,S=s,F=f\right)$
+$-P\left(\hat{Y}\,|\,S=s^{\prime},F=f\right)\bigr{|}\enspace/\enspace ACC_{F=f},$ (3)
+where $ACC_{F=f}$ calculates the accuracy of given different forgery approaches.
+After applying Eq.
+3 to each forgery method, we calculate their results and then obtain the final fairness score by averaging them.
 We refer to this approach as Utility Regularized Metric.
@@
 In summary, we recommend not relying on a single fairness metric but rather considering a combination of multiple metrics to collectively reflect the fairness of a detector.
-5 Results and Analysis
-5.1 Experimental Setup
+Bias Pruning with Fair Activations
+In this section, we present the Bias Pruning with Fair Activations (BPFA) approach, which develops a novel pruning metric combining weights and the fairness of activations to determine weight importance.
+Then, we prune those with the lowest pruning scores based on a predefined pruning rate by layer.
+Here we utilize an unstructured pruning strategy.
+Noting that the proposed BPFA can be directly extended to process other biases except for racial bias.
+Pruning Metric.
+Consider a convolutional layer weights $W$ of shape $(C_{out},C_{in},S_{Ker}^{h},S_{Ker}^{w})$, where $C_{out}$ represents the number of output filters, and each filter has dimension $(C_{in},S_{Ker}^{h},S_{Ker}^{w})$.
+For one data sample, the output of this layer is denoted as $X$ with shape $(C_{out},S_{out}^{h},S_{out}^{w})$.
+The L2 norm by filter of $X$ is represented as $\lVert X\rVert_{2}\in\mathbb{R}^{C_{out}}$.
+We compute the average L2 norm across all samples from a specific race ${s\in\mathbb{S}}$, denoted by $Z^{s}=\lVert X\rVert_{2}^{s}$.
+For each filter, we then calculate the standard deviation of these norms across all races, which serves as the bias for that filter:
+$BIAS_{i}=std\left(\left\{Z_{i}^{s}\right\}_{s\in\mathbb{S}}\right),$ (4)
+where $std(\cdot)$ denotes the standard deviation.
+This bias measures the variability of outputs across different races.
+The pruning score (PS) for each weight $W_{ijkm}$ in the convolutional layer at the position ${(i,j,k,m)}$ is then calculated by combining the weight with respect to the computed bias:
+$PS_{ijkm}=\frac{\left|W_{ijkm}\right|}{BIAS_{i}}.$ (5)
+By comparing the pruning scores, we can identify and potentially remove weights that have the least impact on model utility but the greatest impact on bias.
+This allows us to reduce bias and improve fairness without sacrificing performance.
+Note that while our method is illustrated using convolutional layers as an example, it can be easily extended to linear layers.
+Benchmark Experiments
+Experimental Setup
 Dataset.
@@
 Algorithms.
-We summarize the face forgery detection algorithms in Sec A.4.
+We summarize the face forgery detection algorithms in the section "Face Forgery Detection Algorithms Categories" in Supp.
 For a comprehensive and fair analysis, we select several representative algorithms.
@@
 For frequency-based detectors, we select F3Net [12], SPSL [13] and SRM [14].
+For fairness-enhanced detectors, we select DAG [37](Xception as base model), DAW [37](Xception as base model) and PFGDFD [38](UCF as base model).
 In detail, these models are trained with the Adam optimization algorithm with a learning rate of 0.0002 and an epoch number of 10.
@@
 And data augmentation methods including image compression, horizontal flip and rotation are applied.
-Fairness Metric Spatial-based Frequency-based
-Xception RECCE UCF Capsule FFD CORE F3Net SPSL SRM
-Naive Metric DPD 0.1538 0.1317 0.1782 0.1428 0.1325 0.0914 0.0764 0.0505 0.1335
-DEOdds 0.1672 0.1379 0.1655 0.1267 0.1226 0.0858 0.0877 0.0480 0.0982
-DEO 0.2095 0.1777 0.2334 0.1849 0.1684 0.1200 0.1030 0.0567 0.1625
-STD 0.0563 0.0473 0.0635 0.0513 0.0470 0.0351 0.0278 0.0211 0.0491
-Approach Averaged Metric AADPD 0.1957 0.1644 0.2107 0.1655 0.1579 0.1136 0.1102 0.0592 0.1415
-AADEOdds 0.1674 0.1378 0.1655 0.1267 0.1255 0.0887 0.0951 0.0512 0.0984
-AADEO 0.2099 0.1777 0.2334 0.1849 0.1742 0.1260 0.1178 0.0633 0.1630
-AASTD 0.0721 0.0603 0.0773 0.0613 0.0577 0.0442 0.0425 0.0236 0.0523
-Utility Regularized Metric URDPD 0.1314 0.1158 0.1433 0.1013 0.1047 0.0805 0.0769 0.0342 0.0917
-URDEOdds 0.1069 0.0908 0.1069 0.0782 0.0802 0.0598 0.0624 0.0315 0.0630
-URDEO 0.1437 0.1283 0.1614 0.1128 0.1169 0.0909 0.0842 0.0356 0.1060
-URSTD 0.0482 0.0423 0.0523 0.0374 0.0381 0.0312 0.0295 0.0136 0.0338
-Table 3: Bias evaluation on FairFD for nine face forgery detectors using Naive Metrics, Approach Averaged Metrics, Utility Regularized Metrics.
-For each row, we use red to represent the minimum, followed by orange for the second lowest, blue for the third lowest, and dark green for the fourth lowest.
-5.2 Benchmarking Fairness of Face Forgery Detectors
-The benchmark results (Tab 3) present a comprehensive evaluation of nine face forgery detectors using various fairness metrics.
+However, when applying these data augmentation methods to DAG, we find that its fairness level significantly deteriorated.
+For a fair comparison, we report below the results using data augmentation.
+Meanwhile, the results without data augmentation are presented in the section "Results without Data Augmentation" in Supp.
+Fairness Metric Spatial-based Frequency-based Fairness-enhanced
+Xception RECCE UCF Capsule FFD CORE F3Net SPSL SRM DAG DAW PFGDFD
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2: Bias evaluation on FairFD for 12 face forgery detectors using Naive Metrics, Approach Averaged Metrics, Utility Regularized Metrics.
+For each row, the best values are underlined and bolded, followed by the second-best values which are underlined, bolded, and italicized, the third-best values are bolded, and the fourth-best values are bolded and italicized.
+Figure 3: Detailed utility (AUC) for diverse races, forgery approaches, detectors.
+Benchmarking Fairness of Face Forgery Detectors
+[…]
+Method DPD DEOdds DEO STD
+MA
+DPD
+MA
+DEOdds
+MA
+DEO
+MA
+STD
+UR
+DPD
+UR
+DEOdds
+UR
+DEO
+UR
+STD
+AUC ACC
+   Original 0.0203 0.0304 0.0215 0.0080 0.0556 0.0481 0.0571 0.0219 0.0320 0.0299 0.0324 0.0126 0.6763 0.7618
+   WEIG 0.0183 0.0258 0.0201 0.0072 0.0564 0.0451 0.0586 0.0219 0.0324 0.0277 0.0334 0.0126 0.6769 0.7615
+   RoBA 0.1128 0.1598 0.1395 0.0445 0.1462 0.1616 0.1432 0.0583 0.0893 0.1024 0.0867 0.0356 0.6331 0.7037
+SPSL    BPFA 0.0181 0.0209 0.0200 0.0072 0.0473 0.0357 0.0496 0.0182 0.0265 0.0218 0.0275 0.0102 0.6862 0.8055
+   Original 0.1099 0.1005 0.1242 0.0398 0.1552 0.1196 0.1623 0.0576 0.1037 0.0763 0.1092 0.0384 0.7304 0.5751
+   WEIG 0.1098 0.1003 0.1240 0.0398 0.1550 0.1194 0.1621 0.0576 0.1035 0.0761 0.1090 0.0384 0.7304 0.5751
+   RoBA - - - - - - - - - - - - 0.5967 -
+FFD    BPFA 0.1096 0.0999 0.1237 0.0397 0.1546 0.1189 0.1617 0.0574 0.1032 0.0758 0.1087 0.0382 0.7305 0.5760
+   Original 0.0805 0.1396 0.1032 0.0328 0.1393 0.1560 0.1360 0.0530 0.0881 0.0986 0.0860 0.0335 0.6302 0.6019
+   WEIG 0.0789 0.1340 0.1012 0.0319 0.1349 0.1494 0.1320 0.0513 0.0853 0.0944 0.0835 0.0324 0.6298 0.6021
+   RoBA - - - - - - - - - - - - 0.5468 -
+PFG- DFD    BPFA 0.0594 0.1337 0.0796 0.0238 0.1079 0.1442 0.1006 0.0411 0.0644 0.0969 0.0578 0.0245 0.6445 0.7415
+Table 3: Experiments with different fairness pruning methods.
+We highlight the best method for each metric in bold.
+And we use ’-’ to indicate methods that cause severe performance degradation, rendering the detector unusable even setting a pruning rate as low as 0.1%.
+Benchmark Results.
+The benchmark results (shown in Table 2) present a comprehensive evaluation of 12 face forgery detectors using various fairness metrics.
 We highlight the four fairest detectors using different colors.
-Based on the results, we draw the following significant observations:
-Current face forgery detectors all exhibit a high degree of racial bias.
-Even the DPD of SPSL, which has the smallest racial bias, reaches 0.0505.
-This indicates that the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 5.05%.
-When separately calculating and averaging for each forgery method, the AADPD reaches 5.92%.
-On the other hand, for the least fair detector UCF, the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 17.82%.
-This reminds researchers to address the racial bias in face forgery detection.
-Current face forgery detectors exist racial bias variation.
-Comparing the least fair detector UCF with the most fair detector SPSL, the former’s URDPD is 4.19 times that of the latter, URDEOdds is 3.39 times, URDEO 4.53 times, and URSTD is 3.85 times, showing significant gap in racial bias.
+Based on the results, we draw the following significant observations: (1) Current face forgery detectors all exhibit a high degree of racial bias.
+The DPD metric of SPSL can achieve 0.0203 with the smallest racial bias.
+This indicates that the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 2.03%.
+When separately calculating and averaging for each forgery method, the AADPD metric reaches 5.56%.
+On the other hand, for the least fair detector UCF, the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 17.65%.
+This reminds researchers to address the racial bias in existing face forgery detection models.
+(2) Current face forgery detectors have racial bias variation.
+Comparing the least fair detector UCF with the most fair detector SPSL, the former’s URDPD is 4.76 times that of the latter, URDEOdds is 3.58 times, URDEO 4.98 times, and URSTD is 4.48 times, showing significant gap in racial bias.
 Other detectors also exhibit varying degrees of racial bias.
-Furthermore, we observe that three frequency-based detectors, SPSL, F3Net, and SRM, consistently demonstrate the smallest racial bias across all fairness metrics.
-This indicates that although current detectors are not explicitly designed for racial fairness, detectors based on different mechanisms still possess varying degrees of racial fairness.
-We conduct an in-depth investigation into this in Sec 5.3.
-5.3 Analyses and Discussions
-5.3.1 Balanced Training Dataset.
-Data imbalance across races is a prevalent source of bias.
-To assess the impact of imbalanced training sets on racial bias, we create a dataset that balances across different racial groups.
-The dataset is obtained by sampling from the FF++.
-Because FF++ contains a very few number of African and Indian subjects, we can only select 26 subjects from each race.
-This balanced dataset is employed as the training set.
-We adopt the same data preprocessing methods and training parameters as the racially imbalanced training set.
-Notably, due to the reduced dataset size, the number of epochs is doubled to $20$.
-Metrics Xception F3Net RECCE UCF
-Unbalanced Balanced Unbalanced Balanced Unbalanced Balanced Unbalanced Balanced
-Naive Metric DPD 0.1538 0.0852 0.0764 0.0817 0.1317 0.0836 0.1782 0.0908
-DEOdds 0.1672 0.0837 0.0877 0.0630 0.1379 0.0663 0.1655 0.0674
-DEO 0.2095 0.1076 0.1030 0.1027 0.1777 0.1027 0.2334 0.1104
-STD 0.0563 0.0334 0.0278 0.0329 0.0473 0.0314 0.0635 0.0355
-Approach Averaged Metric AADPD 0.1957 0.1030 0.1102 0.0996 0.1644 0.1044 0.2107 0.0961
-AADEOdds 0.1674 0.0857 0.0951 0.0691 0.1378 0.0746 0.1655 0.0674
-AADEO 0.2099 0.1116 0.1178 0.1149 0.1777 0.1193 0.2334 0.1104
-AASTD 0.0721 0.0412 0.0425 0.0393 0.0603 0.0405 0.0773 0.0374
-Utility Regularized Metric URDPD 0.1314 0.0741 0.0769 0.0770 0.1158 0.0733 0.1433 0.0723
-URDEOdds 0.1069 0.0573 0.0624 0.0511 0.0908 0.0505 0.1069 0.0487
-URDEO 0.1437 0.0824 0.0842 0.0900 0.1283 0.0847 0.1614 0.0842
-URSTD 0.0482 0.0297 0.0295 0.0302 0.0423 0.0284 0.0523 0.0281
-Utility Accuracy 0.5552 0.3984 0.5266 0.3308 0.5011 0.4276 0.5557 0.3682
-Table 4: Evaluations on the unbalanced and balanced training dataset.
-Tab 4 presents the results.
-For convenience, we put the results of models trained on unbalanced and balanced datasets together.
-For both Naive Metric and Approach Averaged Metric, we observe that models trained on balanced datasets generally exhibit lower racial bias.
-Despite the significant improvement brought by a balanced dataset, these detectors still demonstrate a relatively high racial bias.
-It is crucial to note that, particularly for F3Net, not all metrics demonstrate improvement.
-These results indicate that training models on a balanced dataset cannot completely address the issue of racial bias, because there are other factors contributing to racial bias.
-For the Utility Regularized Metric, we note that the enhancement achieved through a balanced training dataset is less significant or may even diminish compared to the Naive Metric and Approach Averaged Metric.
-This is because this set of metrics is influenced by performance, and the reduced data volume results in a notable model performance drop.
-This highlights the advantage of the Utility Regularized Metric, i.e., it can reflect the model’s performance.
-Based on the aforementioned observations, we argue that utilizing a race-balanced training dataset may not be a recommended way.
-The fact that deliberately collecting such balanced data in real-world scenarios usually implies discarding a wealth of available imbalanced datasets.
-Existing datasets are also seldom racially balanced.
-Furthermore, employing a racially balanced training set does not guarantee effective mitigation of racial bias.
-In conclusion, training with race-balanced datasets poses challenges, considering the scarcity of such datasets in real-world scenarios and the limited effectiveness in addressing racial bias.
-Therefore, it is preferable to utilize some other fair learning methods that exhibit better trade-offs.
-Figure 4: Probability score distribution of each race.
-Xception Utility(Accuracy) ↑ Fairness ↓
-Caucasian Asian African Indian STD AccGap
-Best Threshold 0.6530 0.7300 0.5380 0.6660 - -
-Real Face 0.7960/0.8964 0.7201/0.8813 0.8449/0.8767 0.7715/0.8888 0.0450/0.0075 0.1248/0.0197
-FaceSwap 0.8910/0.8265 0.8949/0.7883 0.7668/0.7373 0.8968/0.8264 0.0552/0.0366 0.1300/0.0892
-(a) Performance and fairness with 0.5 and optimal thresholds as threshold respectively. Left of ’/’ 0.5. Right of ’/’ is optimal threshold.
-Threshold AADPD AADEOdds AADEO AASTD URDPD URDEOdds URDEO URSTD
-0.5 0.1274 0.1274 0.1300 0.0501 0.1551 0.1551 0.1507 0.0607
-BEST [36] 0.0544 0.0544 0.0892 0.0220 0.0301 0.0301 0.0497 0.0122
-(b) Fairness metric results at different thresholds.
-Table 5: Threshold Analysis: performance and fairness at different thresholds.
-5.3.2 Setting Different Threshold for Each Race.
-The study in [36] suggests that in face recognition, the confidence score distributions vary among different race groups.
-So, they propose setting different thresholds for different races can enhance fairness as well as improve overall performance.
-We utilize only the real face and FaceSwap portions of FairFD.
-Firstly, we plot the probability score distribution for each race in Fig 4, revealing distinct differences in the confidence score distribution across different races.
-The Asian subset shows a higher frequency of high confidence scores, making it more prone to be classified as a fake face.
-Therefore, a larger threshold can be set for the Asian subset.
-On the other hand, the African subset exhibits a lower frequency of high probability scores, making it less likely to be classified as a fake face, allowing for a smaller threshold.
-Next, we set the optimal thresholds, which are values that maximize the overall performance for each racial subset.
-Tab 5(a) presents the optimal threshold values, as well as the performance and fairness scores when using threshold 0.5 and the optimal thresholds.
-We observe that, under the condition of maximizing overall performance, there is a certain degree of reduction in racial bias.
-Tab 5(b) demonstrates racial bias using a fairness metric, showing that indeed, setting different thresholds can reduce racial bias.
-This conclusion aligns with the findings in [36] consistently.
-5.3.3 Analysis of Frequency-Based Detector.
-Fairness Metric Xception F3Net RECCE UCF
-RGB Grayscale RGB Grayscale RGB Grayscale RGB Grayscale
-Naive Metric DPD 0.1538 0.1309 0.0764 0.0780 0.1317 0.0811 0.1782 0.1324
-DEOdds 0.1672 0.1439 0.0877 0.0976 0.1379 0.0840 0.1655 0.1512
-DEO 0.2095 0.1789 0.1030 0.1035 0.1777 0.1076 0.2334 0.1828
-STD 0.0563 0.0470 0.0278 0.0283 0.0473 0.0295 0.0635 0.0475
-Approach Averaged Metric AADPD 0.1957 0.1673 0.1102 0.1147 0.1644 0.1097 0.2107 0.1722
-AADEOdds 0.1674 0.1439 0.0951 0.1055 0.1378 0.0900 0.1655 0.1512
-AADEO 0.2099 0.1789 0.1178 0.1193 0.1777 0.1195 0.2334 0.1828
-AASTD 0.0721 0.0614 0.0425 0.0441 0.0603 0.0410 0.0773 0.0658
-Utility Regularized Metric URDPD 0.1314 0.1144 0.0769 0.0802 0.1158 0.0785 0.1433 0.1209
-URDEOdds 0.1069 0.0935 0.0624 0.0687 0.0908 0.0603 0.1069 0.0986
-URDEO 0.1437 0.1249 0.0842 0.0859 0.1283 0.0876 0.1614 0.1321
-URSTD 0.0482 0.0419 0.0295 0.0309 0.0423 0.0294 0.0523 0.0461
-Table 6: Comparison of fairness on RGB and Grayscale images.
-In our previous findings, we observe that frequency-based methods exhibit lower racial bias compared to spatial-based methods.
-In this section, we provide a preliminary discussion.
-Spatial-based methods focus on learning forgery clues in the spatial domain, mainly including color mismatch, textures, shapes, and blending boundaries [42].
-On the other hand, frequency-based methods concentrate on learning forgery clues in the frequency domain, especially targeting high-frequency information related to blending boundaries, edges, and textures [12].
-Frequency-based methods capture less color information, which is also crucial in distinguishing between different racial groups.
-Consequently, we make the assumption that frequency-based methods’ racial biases are smaller due to these detectors learning less color information.
-Therefore, we convert the RGB images of our FairFD dataset into grayscale to eliminate color information but retain frequency domain information, and test multiple detectors.
-The results in Tab 6 demonstrate that the racial bias of spatial-based detectors decreases, while the racial bias of frequency-based methods does not decrease and the changes are small.
-Thus, color indeed appears to be a contributing factor that leads spatial-based methods to have higher racial bias compared to frequency-based methods.
-There is still much exploration to be done in the comparative study of frequency and spatial domains, which can significantly contribute to the development of methods for mitigating racial bias.
-We leave this avenue of research for future investigation.
-6 Conclusion
-This paper explores how to evaluate the racial fairness of face forgery detectors.
-We identify numerous disadvantages in existing datasets and fairness metrics, then propose a novel dataset FairFD and two sets of fairness metrics to address these issues.
+Furthermore, we observe that three frequency-based detectors, SPSL, F3Net, and SRM, consistently demonstrate a smaller racial bias across all fairness metrics.
+We conduct an in-depth investigation into this in the section "Analyses and Discussions" in Supp.
+Detailed Utility Results.
+To present more detailed results, we present the AUC for each detector, each race, and each forgery method in Figure 3.
+Results show that different forgery methods exhibit varying levels of utility.
+This validates the advantage of Utility Regularized Metric.
+Detailed Fairness Results.
+We present the standard deviation (STD) of the ACC for the four races in Figure 4 for each forgery method(including Real Face).
+Our findings reveal that different forgery methods exhibit varying levels of fairness, and different detectors rank the fairness of these forgery methods differently.
+This validates the advantage of the proposed Approach Averaged Metric.
+Figure 4: Fairness (STD) for different forgery approaches.
+Evaluating BPFA
+Baseline Algorithm
+We select two baseline methods for comparison.
+The first WEIG uses only the absolute values of the weights as the pruning score, and the second RoBA uses only the reciprocal of the bias of the activations.
+More details about these baselines are shown in Supp.
+For all three methods, we prune the parameters with the lowest pruning scores.
+These pruning baselines can be directly applied in existing SOTA forgery detection models.
+Results Analysis
+The experimental results under optimal pruning rates for each detector and method are shown in Table 3.
+It can be found that the proposed method BPFA consistently outperforms all baseline methods, achieving superior fairness without compromising utility.
+Although WEIG generally preserves good utility and enhances fairness, its improvements in fairness are not as significant as those achieved by BPFA.
+On the other hand, RoBA exhibits highly unstable performance.
+It results in improving fairness but at the cost of significantly reduced utility, which can render the model nearly unusable in the forgery detection task.
+These results prove the superior performance of BPFA in enhancing both utility and fairness for forgery detection.
+It is encouraging to find that SPSL+BPFA achieves the new state-of-the-art performance.
+The only parameter in our method is the pruning rate.
+The ablation study for pruning rate is detailed in three tables in the section "Ablation Study on Pruning Rate" in Supp.
+Our findings indicate that different detectors and different methods have significantly different optimal pruning rates (which correspond to the least decrease in utility (or even improvement) while achieving the best average value across the 12 fairness metrics.
+Analyses and Discussions
+Here we conduct deeper analyses and give some insights: (1) We train detectors using balanced training data, finding that while racial bias can be reduced, the cost of collecting balanced data is substantial; (2) We set the optimal classification threshold for each race and then use the resulting accuracy values to calculate fairness.
+The conclusions show that this method is highly cost-effective and significantly improves both utility and fairness simultaneously; (3) We analyze that frequency-based detectors exhibit superior fairness performance because of not utilizing race-sensitive information, e.g.
+color.
+More analyses and details are shown in Supp.
+Conclusion
+This paper early explores a comprehensive racial bias evaluation benchmark for forgery detection, which provides a newly self-construct dataset, fairness metric and unified protocols.
+We identify numerous disadvantages in existing datasets and fairness metrics, then propose a novel dataset FairFD dataset and two sets of fairness metrics to address these mentioned issues.
+Besides, we also propose a novel Bias Pruning with Fair Activations algorithm to improve the fairness performance without an extra training process.
 Emphatically, we evaluate the fairness of multiple existing face forgery detectors.
-The results indicate varying levels of racial bias in current detectors.
-Further analyses reveal some insights into the emergence of racial bias.
-In summary, addressing fairness is a crucial progress for the advancement of face forgery detection.
-We propose several unsolved problems here and hope that our work will help to make progress towards solving them.
-1).
-Although we provide an image-level evaluation dataset, video-level evaluation datasets would still be important.
-This is because some forgery methods are only applicable at the video level [43, 44, 45], and there are already some detectors based on videos [46, 30] that also suffer from racial bias issues.
-2).
-Although we propose two sets of novel fairness metrics to address the limitations of existing metrics, developing a unified and theoretic metric is still an interesting research direction.
-3).
-Developing fair face forgery detectors that have low racial bias is urgently needed.
-Appendix A Appendix
-A.1 Classification of Forgery Approaches
-There are several approaches for creating fake faces. Here, we provide a brief overview of the classification of forgery methods:
-1.
-Identity-replaced Forgery Approach refers to substituting the original identity in an image, e.g., FaceSwap.
-FaceSwap [21, 22, 47, 48] involves replacing the face of a person in a video or image with another person’s face.
-The method usually uses deep learning algorithms to detect and extract the faces of the two people and then swaps them.
-2.
-Identity-remained Forgery Approach retains the original identity and alters other facial attributes, e.g., Attribute Manipulation and Expression Reenactment.
-Attribute Manipulation [49, 50, 25, 51, 52] involves manipulating the attributes of a face, such as skin color, gender, and eye shape, while preserving the identity of the person.
-The method usually adopts a GAN.
-Expression Reenactment [53, 45, 54, 23, 24] involves transferring the facial expressions of one person to another person’s face.
-The method usually detect and track the facial landmarks of the two people and then transfers the expression from the source face to the target face.
-A.2 Existing Fairness Metrics
-Due to the complexity of fairness evaluation, numerous fairness metrics are proposed from various perspectives to cater to different scenarios.
-Therefore, we need to leverage multiple fairness metrics to evaluate the fairness in face forgery detection comprehensively.
-Here, we present the four most commonly used fairness metrics and outline their respective applicable scenarios.
-Demographic Parity Difference (DPD): In the context of face forgery detection, where label 1 represents fake face, DPD reflects the model’s inclination to categorize faces of a specific race as fake.
-When people perceive the classification of faces as fake as a form of discrimination, we can leverage DPD to assess the extent of bias in the model.
-$DPD=\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}\,\,|\,\,S=s\right)-P\left(\hat{Y}\,\,|\,\,S=s^{\prime}\right)\right|,$ (4)
-where $\hat{Y}$ is the predicted labels. $\mathbb{S}$ represents the set of sensitive attributes, $S\in\mathbb{S}$ and $\mathbb{S}=\{\text{Caucasian},\text{Asian},\text{Indian},\text{African}\}$.
-Difference in Equalized Odds (DEOdds): DPD may fail in certain situations [55].
-Considering a scenario where a face forgery detector classifies faces of African and Caucasian individuals as fake at a similar rate, but the model makes different types of errors for the two groups.
-Specifically, for African faces, the false positive rate is significantly higher than for Caucasian faces, while for Caucasian individuals, the false negative rate is higher.
-In such a case, we can use DEOdds.
-$DEOdds=\frac{1}{2}\underset{y=\left\{0,1\right\}}{\Sigma}\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}|Y=y,S=s\right)-P\left(\hat{Y}|Y=y,S=s^{\prime}\right)\right|.$ (5)
-Equal Opportunity (DEO): DEO has more relaxed conditions compared to DEOdds.
-When assessing the fairness between group A and group B, only the images of faces that are inherently fake need to be considered.
-DEO solely focuses on true positive rates and does not capture the overall classification differences.
-$DEO=\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}|Y=1,S=s\right)-P\left(\hat{Y}|Y=1,S=s^{\prime}\right)\right|.$ (6)
-Standard Deviation (STD): STD differs from metrics mentioned above.
-STD considers the overall variability rather than just the differences between the best and worst-performing ethnicities.
-We use accuracy as the performance metric.
-$STD=std\left(\left\{acc\left(S=s\right)_{s\in\mathbb{S}}\right\}\right),$ (7)
-where $std$ is a function that calculates the standard deviation of given list. $acc$ calculates the accuracy of a given race.
-A.3 Face Forgery Detection Datasets
-We summarize the information of existing datasets in Tab 2. Below we give simple description of some widely used face forgery datasets.
-FaceForensics++(FF++) [1] is a forensics dataset that consists of 1000 original video sequences downloaded from the Internet(,i.e.,YouTube).
-The videos are manipulated with four automated face manipulation methods.
-For face swap, they use Deepfakes [56] and FaceSwap [21].
-For expression reenactment, they use Face2Face [45] and NeuralTextures [57].
-UADFV [26] contains videos of varying classes, with each video being classified as either real or fake.
-The dataset is relatively small, with only 98 videos, but it has been found to be convenient in terms of how the data is formatted.
-CelebDF-v2 [27] is a comprehensive collection for deepfake forensics, comprising 590 real videos and 5,639 DeepFake videos featuring celebrities with high-quality.
-These videos were generated through an enhanced synthesis process, ensuring superior visual quality and better representing the DeepFake content prevalent on the internet.
-Deepfake Detection Challenge(DFDC) [28] is created by Facebook in partnership with other industry leaders and academic experts.
-The dataset consists of 128,154 videos featuring 960 paid actors.
-The dataset was used in a Kaggle competition to create new and better models to detect manipulated media.
-DeeperForensics-1.0(DF-1.0) [29] contains 60,000 videos and 17.6 million frames with 100 consented actors.
-The actors in DF-1.0 have four skin tones: white, black, yellow, brown, with roughly balanced ratio.
-Different from other datasets, DF-1.0 attach importance to high-quality and diversity of source face videos.
-Each actor has various poses, expressions, and illuminations.
-And DF-1.0 has more real videos than fake videos with a ratio of 5:1.
-ForgeryNet [30] is a very large dataset for real-world face forgery detection, with 2.9 million images, 221,247 videos and 15 forgery approaches.
-And a pipeline of conducting various face forgery approaches are proposed.
-A.4 Face Forgery Detection Algorithms Categories
-Current face forgery detectors can be roughly divided into two categories: Spatial-based, and Frequency-based.
-Spatial-based method [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] is based on the spatial domain features(,i.e., forgery clues) of the fake face.
-Such as RECCE [10] utilizes an encoder to reconstruct real face images, thereby exploring the differences between real and fake images in the spatial domain.
-The differences in the reconstruction of the two are then used as guidance to train a classifier for detecting deepfakes.
-Frequency-based method [12, 13, 14] identifies distinctions between real and fake faces in frequency domain, which are used to detect whether the face is forged.
-Such as F3Net [12] utilizes frequency-aware decomposed image components and local frequency statistics to explore forgery patterns, enabling effective forgery detection.
-A.5 Fairness in Face Recognition
-An increasing number of researchers are are now paying attention to societal issues of artificial intelligence, including adversarial example [58, 59, 60], privacy protection [61, 62, 63], and fairness concerns.
-The research on fairness in face forgery detection is similar to the study of fairness in face recognition.
-RFW [32] is first introduced as a race balanced test dataset for face recognition.
-They also propose IMAN which uses a deep information maximization adaptation network to align global distribution to decrease race gap at domain-level, and learns the discriminative target representations at cluster level.
-Another balanced face recognition dataset BFW [36] is proposed to evaluate the fairness of face recognition system both for gender and ethnic groups and this work also shows variations in the optimal scoring threshold for face-pairs across different subgroups.
-The issue of racial bias in machine learning has garnered significant attention in multi fields.
-The research on fairness in face forgery detection is akin to the study of fairness in face recognition.
-RFW [32] is first introduced as a balanced test dataset for ethnic group.
-IMAN [32] uses a deep information maximization adaptation network to align global distribution to decrease race gap at domain-level, and learns the discriminative target representations at cluster level.
-Another balanced faces dataset BFW [36] is proposed to evaluate the fairness of face recognition system both for gender and ethnic groups and this work also shows variations in the optimal scoring threshold for face-pairs across different subgroups.
-DebFace [37] uses a de-biasing adversarial network to extract disentangled feature representations for both unbiased face recognition and demographics estimation and adopts adversarial learning to minimize correlation among feature factors so as to abate bias influence.
-[38] involves multiple preprocessing methods to improve the dual-shot face detector, data re-sampling to balance the data distribution, and multiple data enhancement methods to increase accuracy performance and proposes a linear-combination strategy is adopted to benefit from multi-model fusion.
-Meta Balanced Network [64] uses meta-learning algorithm to learn adaptive margins in large margin loss to mitigate the algorithmic bias in face recognition models.
-MixFairFace [39] proposes MixFair Adapter to determine and reduce the identity bias of training samples.
-A.6 Number of Subjects is Limited
-We propose that the limited number of subjects makes it challenging to draw meaningful comparisons across groups.
-Caucasian Asian African Indian
-TPR 0.7764 0.7365 0.6289 0.6801
-TNR 0.9502 0.9652 0.9314 0.9646
-Table 7: TPR and TNR of each race.
-Here we do a verification experiment to explain and prove it.
-We utilize the FF++ dataset, retains only the Caucasian, Asian, African, and Indian subsets.
-The remaining dataset consists of a total of 677 subjects.
-For each race, 16 subjects are chosen as the test dataset, and the rest serve as the training dataset (9:1 approximately).
-We train an Xception model on the training dataset for deepfake detection and evaluate its performance on the test dataset.
-We follow the hyperparameters specified in [41] and train the model for 10 epochs.
-Figure 5: TPR and TNR for each subject.
-In Fig 5, we show the TPR and TNR on each subject.
-We can observe significant fluctuations in the model’s performance across each subject.
-The presence of extreme results also indicates that subjects significantly influence the model’s classification.
-Therefore, when the number of subjects is insufficient, it becomes challenging to capture the overall characteristics of a group, making it difficult to draw meaningful comparisons across race groups.
-A.7 Visualization in Feature Space
-The research on fairness in face forgery detection is similar to the study of fairness in face recognition(see Sec A.5 of Supp for more details).
-As stated in [32], one reason for racial bias is that different subsets’ features are totally separate.
-So they take the racial bias as a problem of domain gap and propose IMAN(information maximization adaptation network) to decrease this domain gap.
-We sample 500 samples for each ethnic group and use the well trained Xception model as the detector to obtain features for these samples.
-Then, we plot the t-SNE dimensionality reduction graphs for the feature spaces of the four ethnic subsets in Fig 7.
-Considering that different forgery approaches will cause distinct results, we plot for Real Face, FaceSwap and FaceReen respectively.
-Unlike in [32], we do not observe distinct separation between different subsets at feature level.
-Next, we use the MMD(Maximum Mean Discrepancy) to mathematically calculate the feature distances between the Caucasian subset and other ethnic groups.
-The results in Fig 7 demonstrate that although distances show difference, in comparison to the distances in [32], the distances here are all nearly close to zero.
-These results are because they consider the face recognition task in [32], so the models tends to learn distinctive features for each subject, resulting in significant differences at the feature level.
-In contrast, for the task of face forgery detection, the detector does not exhibit the same tendency, leading to smaller differences in feature level.
-Therefore, attempting to enhance the fairness of face forgery detection from a feature-level perspective may not be feasible.
-Figure 6: T-SNE visualization.
-Figure 7: MMD.
+The results indicate the racial bias in current detectors is generally high and prove the advantages of our proposed BPFA.
+Further analyses reveal some interesting insights into the emergence of racial bias.
+We hope the proposed benchmark can inspire more researchers to develop the field.
+In the future, we will explore a unified fairness metric for diverse biases in more kinds of datasets, and construct the video-level forgery detection datasets for more real applications.
+See social impact in Supp.
+Acknowledgments
+This work was supported in part by the National Natural Science Foundation of China under Grant 62306227, Grant 62276198, Grant U22A2035, Grants U22A2096, Grant 62441601 and Grant 62036007; in part by the Fundamental Research Funds for the Central Universities under Grant ZYTS24142, Grant QTZX23083 and Grant QTZX23042; in part by the Key Research and Development Program of Shaanxi (Program No.
+2023-YBGY-231); in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2022QNRC001; in part by the Guangxi Natural Science Foundation Program under Grant 2021GXNSFDA075011; in part by the Shaanxi Province Core Technology Research and Development Project under grant 2024QY2-GJHX-11; in part by Open Research Project of Key Laboratory of Artificial Intelligence Ministry of Education under Grant AI202401, in part by the Nanning Scientific Research and Technological Development Project 20231042; in part by the ‘111 Center’ (B16037).